### v2

# Colloquium: Synthetic quantum matter in non-standard geometries

Abstract
Quantum simulation is making a significant impact on scientific research. The prevailing tendency of the field is to build quantum simulators that get closer to real-world systems of interest, in particular electronic materials. However, progress in the microscopic design also provides an opportunity for an orthogonal research direction: building quantum many-body systems beyond real-world limitations. This colloquium takes this perspective: Concentrating on synthetic quantum matter in non-standard lattice geometries, such as fractal lattices or quasicrystals, higher-dimensional or curved spaces, it aims at providing a fresh introduction to the field of quantum simulation aligned with recent trends across various quantum simulation platforms, including atomic, photonic, and electronic devices. We also shine light on the novel phenomena which arise from these geometries: Condensed matter physicists may appreciate the variety of different localization properties as well as novel topological phases which are offered by such exotic quantum simulators. But also in the search of quantum models for gravity and cosmology, quantum simulators of curved spaces can provide a useful experimental tool.
Contents
- I Introduction and outline
- II Implementation - engineering exotic geometries
- III Applications - novel phenomena in exotic geometries
- IV Conclusion and Outlook
I Introduction and outline
As envisioned by R. Feynman in 1982 [85], a quantum simulator is a controllable quantum system that is used to imitate another quantum system of interest. Since Feynman’s early inspiration, many quantum simulation platforms have now been developed [7], ranging from atomic to photonic and electronic devices. Common to all these systems is their synthetic nature providing control over the Hamiltonian. Quantum simulators have been used to study quantum phenomena across the whole energy spectrum, from low-temperature behavior of quantum materials to out-of-equilibrium phases and quantum chemistry to high-energy physics and cosmological models. Importantly, the application of quantum simulators has not remained limited to the mere imitation of existing systems, but an increasing amount of experimental and theoretical efforts are dedicated to also implementing quantum systems without real-world counterparts. Through these efforts, the quantum-mechanical consequences of exotic geometries, as an alternative to the Euclidean three-dimensional space we are used to, become the attainable subject of modern science.
Examples of such exotic geometries include various non-periodic lattices, from prototypical quasiperiodic models in one or two dimensions to fractal lattice structures with a fractal Hausdorff dimension, but also curved lattices, for instance, hyperbolic lattices. In addition, the scheme of synthetic dimension allows to build systems with more than three space-like dimensions, and in this way expose quantum systems to an “augmented quantum reality”. The motivations for these efforts are manifold: Space-time curvature is a key element of cosmological or quantum gravity models but typically occurs on scales that are hard to probe experimentally. Quantum simulators of curved spaces bring quantum field theory in curved space-time into the lab and thereby provide a tool for analog studies of cosmological models [278], Unruh effect [280, 217, 142, 14], or analog black holes [98, 192, 250].
Also, on a totally different length scale, in the context of microscopic behavior in condensed matter, synthetically generated settings can provide a fresh perspective on fundamental aspects. Various quantum aspects of matter, both on the single- and the many-body level, strongly depend on the dimensionality and geometry of the space in which matter is embedded. Examples are quantum transport, localization and thermalization behavior, topological transport, and topological classification of quantum systems. For instance, localized 1D or 2D systems with a mobility edge are exclusive to carefully designed quasiperiodic potentials [30, 97, 170, 70, 8]. In synthetic quasicrystals with interactions, the nearly perfect isolation from the environment has enabled the observation of a phenomenon known as many-body localization (MBL) [232, 167], the absence of thermalization. The 4D quantum Hall effect is an exciting example of a topological system that involves topological concepts, known as the second Chern number, beyond the ones relevant to real-world 2D topological matter [145]. Fractal lattices provide a playground to explore topological order without a true bulk [189, 199, 122, 94, 31]. Quasicrystals, characterized by symmetries impossible in regular crystals, open the door to novel symmetry-protected topological phases [273, 82].
In the present colloquium, we showcase implementations of quantum systems in exotic geometries across different experimental platforms and then discuss the novel physical phenomena that can be addressed in these systems. Specifically, the geometries in the focus of this article are:
Quasicrystals.
Quasicrystals are a third kind of solids, between the periodic crystals and the amorphous solids or glasses, which, typically due to a non-negligible amount of disorder, lack translational symmetry. The structure of quasicrystals follows a deterministic construction rule, which gives rise to long-range order as in ordinary crystals. However, like glasses, they are aperiodic. Physically, their order becomes manifest, for instance, through Bragg peaks in a diffraction experiment, which also has played a crucial role in the discovery of quasicrystals in 1984 [237, 154], when D. Shechtman and co-workers found aluminum-manganese alloys with tenfold rotational symmetry [237]. Indeed, some definitions of a quasicrystal demand that it exhibits a rotational symmetry incompatible with periodic structures limited to two-, three-, four-, or sixfold rotational symmetry. According to a more general definition that also admits 1D quasicrystals, this term is simply an abbreviation for quasiperiodic crystal [162].
Shechtman’s discovery, awarded the Nobel Prize in Chemistry in 2011, extended the research on aperiodic tilings from mathematically and/or aesthetically grounded studies in the 1970s [202] to physics and material science. Nowadays, research on quasicrystals is gaining importance through the rapid progress of synthetic materials, which have opened the door to experimental studies of an even larger variety of quasicrystals, including paradigmatic 1D models. On the tight-binding level, these models are of the form
with () creation (annihilation) operators on site . The parameters and are neither periodic (as in crystal models) nor random (as in glassy models), but (at least) one of them is obtained from a quasiperiodic sequence, such as the Fibonacci sequence, see Jagannathan [126] for a review. For instance, the diagonal Fibonacci quasicrystal has constant hopping, , but the potential is a sequence of binary potential values. This sequence is generated iteratively by merging the previous two iterations. Another famous 1D quasicrystal model is the Aubry-André (AA) model with
Here, is irrational, and a random phase. We provide an overview of different realizations of synthetic quasicrystals in the first row of Table 1.
Fractals.
Another non-periodic geometry are fractals. Following a deterministic construction scheme, they may appear rather artificial, but fractals are omnipresent in organic nature [176]. Fractals are characterized by self-similarity: repeating the same patterns on different scales and by a fractal dimension, a non-integer Hausdorff dimension [176]. The concept of Hausdorff dimension is based on box-counting: the method considers a grid with spacing of the embedding Euclidean space and counts the number of boxes that fit into this space. The Hausdorff dimension is determined by the scaling of in the limit , specifically
An example of a fractal, the so-called Sierpiński triangle or gasket, is shown on the top of the second column of Table 1. The structure is generated by dividing an equilateral triangle into smaller ones and removing the central smaller triangle. The Hausdorff dimension of this structure is . A similar structure, starting from a square instead of a triangle, is known as Sierpiński carpet and has a Hausdorff dimension of . The procedure can also be applied to structures embedded in a three-dimensional space, e.g., tetrahedrons. By division into smaller and smaller tetrahedrons combined with the removal of selected tetrahedrons, one ends up at the Sierpiński tetrahedron, which curiously has an integer Hausdorff dimension of 2. Through synthetic systems, fractals have also entered the arena of quantum physics, see Table 1, and in particular, the dynamics of quantum particles in fractal space have become the subject of intense research.
Lattices with extra dimension.
Even integer dimensions can be considered quite exotic if the number of dimensions exceeds three. Edwin Abbott’s satirical novella Flatland from 1884 describes a society of flatlanders living in a 2D world [1]. For them, a 3D sphere that crosses flatland is perceived as a circle that, strangely enough, appears out of nothing and constantly changes its radius until it again shrinks to a point and disappears. The only way to understand the origin of this strange object is by leaving flatland. Modern quantum simulators allow us to leave our 3D perspective on the world, by engineering quantum systems that incorporate other than purely spatial degrees of freedom in a space-like fashion, see Table 1. These experimental systems have opened a door to phenomena occurring only in higher-dimensional spaces, particularly novel types of topological matter [208].
Curved lattices.
Finally, lattice systems can become exotic if their metric is not the usual Euclidean but exhibits some curvature. Curved spaces have been traditionally investigated in high-energy physics and cosmology. In recent years, there has been a growing interest in realizing tabletop simulations of hyperbolic lattices. The motivations arise from discovering the holographic principle [174, 299] and the characterization of classical and quantum states in spaces with negative curvature. One possible way to investigate curved lattices beyond the Euclidean case is to work on the Poincaré disk. Here, the tessellation of the space is defined via two indices , known as the Schläfli symbol. It corresponds to a tessellation of the plane by regular -gons with coordination number [62]. Given a , the curvature of the space is defined by
If , the space is spherical (positive curvature); whereas, if , the space is Euclidean (zero curvature), and finally, if , we have a hyperbolic space (negative curvature) [38]. Euclidean tesselations are limited to the well-known triangular , square , and hexagonal lattices. The example of a hyperbolic lattice is shown in the illustration of Table 1 (right column). Hence, by the choice of connectivity in a lattice , non-Euclidean spaces can be implemented, and in this way, hyperbolic lattices, characterized by negative curvature, have been produced in photonic and electronic quantum simulators, see Table 1. Another approach to curved spaces is through an analogy between sound propagation on a background hydrodynamic flow and field propagation in curved space-time, see Unruh [270]. This analogy has allowed for experimental studies of gravitational phenomena in flat quantum simulators.
This colloquium reviews how the described variety of exotic geometries is implemented and exploited in contemporary quantum simulators using atomic, photonic, and electronic platforms. Specifically, in the next section (Implementation, Sec. II), we will review state-of-the-art quantum engineering techniques used to achieve such geometries. In the subsequent section (Applications, Sec. III), we will discuss the specific implications of exotic geometries on physical system behavior. Here, we will differentiate between three different phenomenological areas: (i) localization phenomena, (ii) topological phenomena, and (iii) analog black holes and cosmology.
II Implementation - engineering exotic geometries
We will present an overview of state-of-the-art quantum simulation systems/techniques by demonstrating how exotic geometries are engineered in different platforms, covering atomic, photonic, and electronic systems as outlined in Table 1.
II.1 Atomic systems
Ultracold atoms are a formidable experimental platform for quantum simulation due to their large tunability of confining potentials and interaction strength and ability to reach strongly-correlated quantum states with long lifetimes. Neutral atoms are routinely cooled to Mikrokelvin temperatures via laser cooling and further cooled to Nanokelvin temperatures via evaporative cooling in magnetic or optical traps. Ultrahigh vacuum conditions allow for thermal insulation from the vacuum chamber walls. Atoms exist as bosons or fermions depending on the total spin of the atomic isotope, and at Nanokelvin temperatures, they form Bose-Einstein condensates (BEC) or degenerate Fermi gases, respectively.
Optical lattices formed by the interference of laser beams then provide a means to study solid-state physics in a system with 10,000 times larger lattice spacing than electronic materials. This makes the optical imaging of each particle possible. From the resulting snapshots of many-body wave functions, arbitrary correlation functions can be extracted, extending our understanding of what observables are experimentally accessible.
Atoms have internal hyperfine states that can equally be controlled and harnessed as an effective spin. This allows to study quantum magnetism or to engineer spin-orbit coupling via two-photon Raman transitions. Some atoms also have long-lived optically excited states that can be employed as orbital degrees of freedom accessed by single-photon transitions.
An important technique for further tuning synthetic systems is Floquet engineering [40]. The Floquet theorem states that a periodically-driven system has an effective description via a static Floquet Hamiltonian when probed stroboscopically. This effective Hamiltonian can have new desired properties, such as stable trapping of a rotating saddle point used in ion traps or artificial gauge fields for neutral particles such as cold atoms or photons [77]. Remarkably, Floquet driving is also a way to go beyond systems with a static counterpart, as in anomalous Floquet topological insulators that break the bulk-boundary correspondence of static systems [222]. Floquet engineering is used in all platforms considered here, i.e., ultracold atoms [77, 293], photonic systems [197] and solid-state systems [195, 221].
With this large toolkit available, ultracold atoms have been used to study diverse effects of solid-state physics, high-energy physics, and quantum systems without counterparts in the real world, as further discussed in the following.
II.1.1 Quasicrystals and fractals
Quasicrystal optical lattices.
Optical lattices have been a primary tool for quantum simulation with cold atoms, including exotic geometries [298]. Very elegantly, optical lattices can also produce quasicrystalline potentials by an appropriate choice of laser beams forming the lattice. The AA model in Eq. (2), a 1D lattice with quasi-random on-site energies, was realized by adding a second weaker 1D optical lattice of incommensurate spatial period due to an incommensurate wavelength [216, 213]. This approach can be used as an alternative to laser speckle disorder [32], but as will be discussed in Sec. III.1 also allows for fine-tuned localization properties, such as 1D mobility edges. In 2D, fivefold quasicrystal potentials would be formed by the interference of five lattice beams [227], and a challenge would be to control the relative phases of the laser beams [61]. This technical challenge can be overcome using the multi-frequency lattice scheme leading to a pairwise interference of the beams at different frequencies with well-controlled relative phases [141]. An eightfold quasicrystal potential was realized by superposing four 1D optical lattices at slightly different frequencies such that they do not interfere with each other [276, 228] — see Figs. 1a,b. These experiments have also triggered new theoretical interest in many-body physics in quasicrystal potentials and new descriptions such as Hubbard models for quasicrystalline potentials [128, 100, 103, 317].
Optical tweezers.
Atomic systems can also be assembled one by one via optical tweezers — see Fig. 1c. These are tightly focused optical dipole traps directly loaded from a magneto-optical trap, where the collisional-blockade mechanism ensures occupation with no more than one atom [229]. Subsequent detection and rearrangement of the traps then allows for the creation of defect-free arrays [83, 20], which are now widely used as starting point for quantum computing protocols and quantum simulation [234]. This approach allows creating arbitrary geometries including Fibonacci quasicrystals [287] and fractal structures [264] — see Fig. 1d. Due to the micrometer-spacing of the tweezers, significant tunnel coupling between them was only reached for the lightest atom lithium [247], but Rydberg excitations can hop between the tweezers and realize relevant quantum many-body models.
Momentum-space lattices.
As an alternative to optical lattices and optical tweezers, one can also use a mapping between real-space lattices and synthetic lattices composed of discrete internal [34] or external [10, 209] states. The external states can be discrete momentum states coupled by two-photon Bragg transitions of appropriate frequencies — see Figs. 2a,b. Synthetic lattices offer the great advantage of microscopic control over all system parameters, such that local gauge fields, tailored disorder, or quasiperiodic potentials can be engineered [11, 8, 289]. Extensions to 2D momentum-space lattices have been discussed [3].
II.1.2 Lattices with synthetic dimension
The synthetic lattice made up of Raman-coupled internal states can be combined with a real-space lattice, such that the internal states provide an additional compact dimension to the system [34]. This approach facilitates imprinting a Peierls phase on the coupling element within this synthetic dimension, giving rise to an artificial gauge field. Furthermore, the synthetic dimension has a natural sharp edge. Exploiting these features, several groups have realized a synthetic quantum Hall stripe and directly observed the skipping orbits at the system’s edge [175, 251, 50]. Due to the quadratic Zeeman splitting between internal atomic levels labeled by the magnetic quantum number , it is also possible to spectrally select different Raman transitions, and depending on the polarisation of the laser beams, transitions can either couple states with or . This feature has been proposed to realize periodic boundary conditions or lattices with nontrivial connectivity such as tori or Möbius strips [35, 13]. A synthetic topological Hall cylinder was recently realized using four internal states of 87Rb atoms coupled with Raman beams and microwave fields [155]. Finally, using the large spin of 162Dy atoms with 17 magnetic sublevels, a 4D system was recently built by combining two real dimensions with two synthetic dimensions in cylindrical geometry given by the magnetic projection and its remainder , respectively [39]. A very short synthetic dimension of just two spin states can be used to emulate a bilayer system, with nontrivial hopping between the layers induced by Raman transitions. This can realize tunable graphene-Haldane bilayers [54] or systems analogous to twisted bilayer graphene featuring flat bands [226].
II.1.3 Curved space
An optical lattice may also mimic curved space if the hopping amplitudes are modified by the coupling to a gauge field that carries information on the effective curvature [33]. The spatially dependent complex hopping amplitudes can be engineered, e.g., by tailoring the profile of the laser beams that drive the laser-assisted hopping as Raman transitions between the two sublattices of a spin-dependent optical lattice. It is predicted that this allows the simulation of the motion of massless fermions in a Rindler metric.
While inhomogeneous tunnel elements have not been experimentally realized so far, there are experiments in bulk systems utilizing inhomogeneous interactions instead. In Viermann et al. [278], inhomogeneous interaction energy results from an inhomogeneous density distribution in a trapped system and can be used to engineer a metric. In Di Carli et al. [71], an inhomogeneous interaction strength was engineered using a magnetic Feshbach resonance and a magnetic field gradient.
A different approach to simulating curved spacetime with ultracold atoms utilized the equivalence of an accelerated frame to a parametric modulation of the interaction strength of an atomic BEC [117]. The protocol leads to matter-wave jets, which share their characteristics with the Unruh radiation, namely long-range coherence combined with local thermal distribution.
An exciting way of studying an expanding universe is to work with BECs in supersonically expanding toroidal traps [29]. This was realized with 1D ring traps with variable radius created via a digital micromirror device [78, 18], finding a redshift of long-wavelength excitations as in an expanding universe as well as the production of solitons and vortices. A 2D version, a bubble trap forming the surface of a sphere [99], was recently realized via radiofrequency dressing of a magnetic trap using the microgravity conditions on the international space station [47]. The experiment studied the thermodynamics of ultracold bosonic atoms, demonstrating substantial cooling upon increasing the size of the sphere, which links to expansion physics analogue to the expanding universe. Such 2D bubble traps are also relevant for the study of space-curvature effects on quantum many-body physics, e.g., on superfluidity and the dynamics of vortices and solitons [44, 267, 168, 268, 266].
II.1.4 Interactions
A central feature of ultracold atoms is the tunability of the interactions. External fields can enforce resonances between different atomic levels, strongly affecting the atomic scattering length. In particular, magnetic Feshbach resonances have become a standard tool to tune the contact interaction strength continuously [56]. In a quantum simulator, this allows to connect the (usually) understood non-interacting regime to the computationally complex interacting regime.
Next to these contact interactions, there are efforts to engineer long-range interactions, e.g., by using dipolar atoms, molecules, optical cavities, or Floquet engineering. In optical tweezers, typically spaced by a few micrometers, strong interactions can be induced by exciting to high-lying Rydberg levels [123, 295]. Contact interactions generally translate to infinite-range interactions in artificial dimensions. In the case of repulsive bosonic atoms in 1D momentum-space lattices, they result, however, in an effectively attractive, finite-ranged interaction, which can be understood via the extra repulsive exchange energy of a symmetrized two-body wave function for distinguishable momentum states [9]. Therefore, interesting many-body physics, such as solitons, are expected in these systems. In the case of synthetic dimensions formed by internal states, it has been proposed to recover short-range interactions required for fractional quantum Hall (FQH) liquids by spatially separating the different internal spin states using a magnetic field gradient [50]. Interactions also decay with distance in an artificial dimension formed by the levels of a harmonic trap [209].
II.2 Photonic systems
Another powerful quantum simulation approach is due to the confinement, propagation, and emission of photons via tailored light-matter interactions. It significantly benefits from the advancements in nanofabrication technologies, enabling the construction and precise sculpting of materials at the optical wavelength scale. Furthermore, the availability of various optical spectroscopy techniques—such as time-resolved measurements, Fourier transform imaging, and interferometry—paired with state-of-the-art light sources and photodetectors facilitate the injection, probing, and monitoring of photonic states (both amplitude and phase) in real and momentum space.
Polaritons, quasiparticles arising from the strong coupling regime between confined photons and dipole excitations (electrons oscillating at the metal surface for surface plasmon polaritons and excitons in semiconductors for exciton-polariton), are one of the most successful photonic platforms for quantum simulation. Potential landscapes in 1D are simulated by harnessing superlattices of multilayers along the propagation direction [6, 206, 112], or by implementing lateral confinement in photonic microwires through modulating the wire width [255, 192, 15, 101]. An example of such modulation is shown in Fig. 3a. Moreover, simulations of 2D tight-binding Hamiltonians are obtained via lateral coupling between micro pillars [125, 225, 136]. In this scheme, the pillar sizes dictate the on-site energies and the hopping terms given by their inter-distances. A crucial advantage of the modulated microwires and the coupled micropillar configurations is the out-of-plane access of spectrally resolved photons escaping the microstructures to far-field detectors. Indeed, a single snapshot of this emission provides tomography of confined polaritonic states in both real and momentum space, as illustrated in Figs. 3b,c. Moreover, the pseudospin texture and phase pattern of the polaritonic states can be directly mapped by polarization analysis [225] and interferometry measurements [225, 125] applied to the out-of-plane emission, respectively. Finally, the matter component of polaritons enables two distinct interaction schemes: (1) self-interaction, facilitating the exploration of many-body physics in the quantum fluid of light [192], and (2) interaction with external magnetic fields, breaking time-reversal symmetry for applications in topological photonics [136].
Another primary photonic platform for quantum simulations involves arrays of coupled single-mode waveguides. These arrays can be easily fabricated in bulk glass using standard lithography techniques with femtosecond lasers, providing a versatile playground for simulating tight-binding Hamiltonians. This is achieved by engineering the evanescent tunneling of light propagating along a waveguide to its neighbors. Paraxial propagation along the direction of the waveguide is described by , equivalent to the Schrödinger equation if the third spatial dimension is equated with time. The onsite and hopping terms of the Hamiltonian are dictated by the refractive index of the waveguides and the spacing between them, respectively. The detection scheme in waveguide array systems is limited to only photographs of near-field intensity in real space. Still, it provides an elegant scenario for monitoring the time evolution of the injected wave packets by capturing the intensity profile at different positions, thereby studying various transport problems [91, 233, 257, 301]. Most importantly, this platform offers a unique scheme to “dynamically” modulate the Hamiltonian by implementing -dependence in the hopping terms. This paves the way for simulating Floquet physics in topological photonics [214, 31] and realizing topological pumping using synthetic momenta [144, 319], as illustrated in Figs. 4a-d.
II.2.1 Quasicrystals and fractals
Inspired by pioneering works in solid-state physics on Fibonacci superlattices [181, 265], early studies on photonic quasicrystals concentrated on Fibonacci superlattices within the polaritonic platform [6, 206, 112]. The first study by Albuquerque and Cottam [6] introduced 2D electron gas (2DEG) sheets at the interfaces of alternating dielectric layers following the Fibonacci sequence and predicted the formation of bulk bands and surface branches of surface plasmon polaritons. Subsequent research extended this scheme to exciton polaritons in Fibonacci-spaced multiple quantum wells [206, 112, 294]. In particular, the experimental results in Hendrickson et al. [112] highlighted the quasicrystal’s long-range order, resulting in the formation of a polariton bandgap, and its lack of periodicity, which promotes efficient out-of-plane light emission extraction.
Contemporary research on quasicrystals in polaritonic platforms has predominantly employed the modulated microwire geometry to study topological and critical properties of chains with Fibonacci quasiperiodic potential: The opening of minigaps obeying the gap labeling theorem (see also Sec. III.2.1) and the log-periodic oscillations of the integrated density of states have been observed in Tanese et al. [255]. Subsequent work has used localized states at the interface of two Fibonacci polaritonic wires to measure topological invariants associated with the gap labeling [15]. Finally, the critical localization behavior of the Fibonacci chain (see also Sec. III.1.1) has been explored in Goblot et al. [101] by continuous interpolation from the AA potential to the Fibonacci, revealing multiple localization/delocalization transitions.
As an alternative to polaritons, waveguide arrays provide another fruitful platform for exploring quasicrystal and fractal physics. Studies of “stationary” localized states in 1D quasicrystals have experimentally revealed the localization phase transition of the AA model [148] (see also Sec. III.1.1). The topological characterization of different 1D quasicrystals via edge states and Chern numbers has been demonstrated in Verbin et al. [275], as well as the adiabatic pumping of localized edge states within 1D quasicrystals by “dynamically” sweeping the phase in the AA model, achieved via slowly tuning the spacing between waveguides in Kraus et al. [144], (see also Sec. III.2.1). Adiabatic pumping provides an analogous behavior to 2D quantum Hall physics within a 1D quasicrystal. Going from 1D to 2D quasicrystals, earl works tackled linear and nonlinear transport and defect dynamics [91]. In this study, quasicrystals were generated by inferential lithography into a birefringent and nonlinear bulk crystal, demonstrating similarities between light propagation and electron tunneling in quasiperiodic potentials. This study also uncovered soliton formation at high intensities and dislocation dynamics when interactions between crystal sites are permitted. Waveguide arrays have also been utilized for the simulation of fractal physics: quantum transport in fractal networks has been explored in Xu et al. [301], suggesting that anomalous transport is governed solely by the fractal dimension, although theoretical studies show that transport properties can become extremely sensitive to the connectivity at bottlenecks of the fractal [218]. Finally, by investigating driven fractal lattices within the framework of Floquet physics, fractal photonic topological insulators have been observed experimentally [31]. Remarkably, this work demonstrated that topologically protected chiral edge states can exist despite the absence of bulk bands.
II.2.2 Lattices with synthetic dimensions
Photonic systems offer a diverse range of degrees of freedom that can be engineered as synthetic dimensions, enabling the study of higher-dimensional physics using platforms with fewer physical dimensions. A straightforward choice for a synthetic dimension in photonics is the polarization state of light. For example, polarization conversion during light propagation in a photonic lattice can be mapped into an effective coupling between two lattices [80]. Other options, based on the nature of photonic modes, include the orbital angular momentum of light [46, 45, 283] or the spatial form of the eigenmodes [171]. In the quantum regime, correlations between photons can be mapped to the synthetic dimension of Fock space. This mechanism can be combined with other degrees of freedom, such as polarization, to achieve more complex mappings [80, 79].
In addition to using properties of photonic states, another strategy to harness synthetic dimensions in photonics is to employ parameters of the Hamiltonian of a photonic lattice that can be precisely modified. For instance, the physics of the 4D quantum Hall effect can be investigated using 2D quasicrystals, where the two modulation phases and act as synthetic momenta [145, 319]. The lattice model and implementation scheme are depicted in Figs. 4a,b, and the experimental results, highlighting corner-to-corner adiabatic pumping, are shown in Fig. 4c.
Interestingly, frequency and time-bin, both related to the temporal aspect of space-time, are two of the most commonly used degrees of freedom for synthetic dimensions in topological photonics. Discrete resonances in ring resonators, shown in Fig. 5a, provide a frequency lattice where nearest-neighbor couplings can be engineered on-demand using electro-optical modulation [198, 310]. This configuration has been implemented in various material platforms, successfully emulating different facets of topological physics [75, 74, 285, 17, 201]. Regarding using time-bin as a synthetic dimension, the scheme involves studying light propagation in two loops of optical fiber of different lengths connected by a beam-splitter. This setup makes the system equivalent to a 2D lattice, where one dimension corresponds to the location of the quantum walk and the other to the time step [297, 296, 291], as shown in Fig. 5(b). Additionally, tuning the beamsplitter behavior over time can introduce another synthetic dimension to explore new topological phases, such as the anomalous Floquet metal [2].
Beyond synthetic coordinates, another essential concept is the synthetic magnetic field (or pseudo-magnetic field). Photons do not interact with magnetic fields directly; therefore, to emulate quantum Hall states in optics, it is necessary to engineer a photonic Hamiltonian that exhibits a synthetic magnetic field. This can be accomplished by deforming a photonic honeycomb lattice, similar to how strained graphene behaves, leading to the experimental observation of Landau levels in coupled micropillar systems [127] and photonic crystal systems [19, 21]. Combining the ultrastrong nonlinearity of single polaritons with a synthetic magnetic field has ultimately enabled the first realization of bosonic analogs of FQH physics [59]. In this setup, single polaritons are formed by the strong coupling between ultra-cold Rydberg atoms and photons in a twisted cavity. This cavity is specifically designed to create a synthetic magnetic field for photons. The observation of angular-momentum-dependent two-photon correlations evidences the creation of Laughlin states.
Synthetic momenta, which differ from synthetic dimensions in spatial space, have recently emerged as a promising concept for exploring higher-dimensional topology. The key idea is to use spatial quantities as synthetic momenta, thereby naturally breaking time-reversal symmetry in the extended momentum space. For instance, in misaligned photonic gratings that share the same period and are evanescently coupled, the relative shifts between the gratings serve as synthetic momenta [191, 150, 190], as shown in Fig. 5c. Using this concept, chiral edge state [191] and 2D nodal lines [150] were proposed for (1+1)-dimensional systems of bilayer gratings, while Weyl semimetal physics with Fermi arc reconstruction [190] was proposed for (1+2)-dimensional systems of trilayer gratings. Recently, synthetic momentum has been employed to explore non-orientable manifolds to circumvent the Nielsen–Ninomiya theorem [87]. Given the vibrant field of multilayer photonics and various recent realizations [259, 258], the experimental implementation of these synthetic momenta systems in the near future is anticipated, potentially addressing 4D or 5D physics.
II.2.3 Curved lattices
The propagation of light in a deformed graphene-like lattice of waveguide arrays has been used in Sheng et al. [238] to simulate a gauge field associated with a topological cosmic string. In this experiment, the strain-induced deformation introduces curvature into the effective Dirac equation, which is then mapped onto the lattice Hamiltonian. The degree of curvature experienced by the photons can be tuned by altering the strain strength within the lattice, and it has been demonstrated that the presence of a cosmic string can localize the propagation of wave packets in comparison to flat space.
II.3 Electronic systems
The last platform we will consider in this section consists of creating artificial lattices in electronic-based systems. We will mainly introduce five different approaches in electronic platforms: (i) electrons on surface, (ii) molecular self-assembly, (iii) twisted layers of two-dimensional materials, (iv) classical electric circuits, and (v) circuit quantum electrodynamics (QED). In these cases, the goal is to simulate a specific lattice model. Some of these platforms are limited to simulate single-particle physics.
In case (i) of electrons on a surface, the electrons of a 2DEG are forced to a specific potential with a specific geometry. Usually, the Shockley surface states of metals are considered for the 2DEG [239], e.g., Cu, or Ag on the (111) surface. The lattice potential is created by placing atoms or molecules with atomic precision using the tip of a scanning tunneling microscope (STM). The tip of the STM is successively used for measuring the single-particle local density of states (LDOS) at the different positions by the differential conductance of the system as a function of the applied tip-substrate bias . This quantity is proportional to the modulus square of the system wave function at a given position and voltage. The major drawback of this technique is the finite lifetime of the surface electronic states. Adatoms of the surface scatter surface electrons into the bulk of the system, leading to the broadening of the electronic resonances. The broadening at 4.6 K is generally smaller in Ag compared to Cu [81]. However, the adatoms placed on Cu, especially CO molecules, are more stable and easier to move on Cu than Ag, making Cu the metallic reference system for creating artificial lattice structures [93]. This combination has been employed for designing lattices with translational invariance [102, 243, 242, 129, 92, 159].
The molecular self-assembly approach deposits molecular precursors on a metallic surface; after annealing, the precursors rearrange themselves in structures that can be more complex than the original precursors, either by following a specific pattern intrinsic to the precursor or because of some property of the substrate. Successively, the structure is analyzed either by STM or by an atomic force microscope [205].
While the before-mentioned approaches are, at present, limited to single-particle physics, twisted two-dimensional materials provide tunable lattices that showcase strongly correlated physics. A prominent example is twisted bilayer graphene exhibiting unconventional superconductivity at a magic twist angle [43]. Quite generically, twisting 2D lattices leads to a Moiré superlattice determined by the twist angle, yielding the basis for tunable electronic materials that can be used for quantum simulation [131]. One advantage of this platform is that adding more layers or combining different materials (e.g. layers of transition metal dichalcogenides) permits the exploration of quantum simulators with diverse properties.
The classical electric circuits approach is based on Kirchhoff’s law in the alternating current regime. The general idea is to represent an electrical circuit by a graph in which the nodes and edges correspond to the circuit junctions and connecting elements. The elements of the circuit can be all linear, such as resistors, capacitors, and inductances. The current conservation can be expressed in a compact form as
where and are vectors of the input electrical current and voltage at all nodes , here, the Laplacian is defined via the matrix of adjacency of the conductances , the list of the total node conductance , and the circuit ground . The impedance to the ground of the node , can be fully determined by with eigenvalues [149].
In addition to capacitors and inductances for the circuit QED approach, the circuit will contain non-linear elements such as Josephson junctions. In this case, it can be mapped into an effective many-body Hamiltonian, including the Jaynes-Cummings Hubbard or simpler hopping models [230]. Circuits of this type operate in microwave frequency regimes.
II.3.1 Quasicrystals and fractals
Electron on surface.
A finite portion of the Penrose tiling has been realized using this technique Collins et al. [60], see Fig. 6a. The various rhombi constituting this quasicrystal have been obtained by placing a CO molecule at the center of each of them, see Fig. 6b. The spectral properties of the system have been analyzed by measuring the LDOS on the various tilings of the quasicrystal. On the different tilings, the LDOS is different, reflecting the distinct nature of the wave function over the quasicrystal.
In Kempkes et al. [130], the electron-on-surface method has been used to implement three generations of the Sierpiński gasket, see Figs. 6c,d. The experiment directly measured the fractal dimension using the box-counting method. For the Hausdorff dimension of the electronic wave function, a value 1.58 was obtained at all energies, compatible with the Hausdorff dimension of the Sierpiński triangle, and in contrast to the value of 2 obtained for electronic wave functions in a square lattice [130]. Additionally, the real space analysis of the wave function for different energies displays a standing wave pattern originating from the interference of the electrons scattered by the CO molecules. The Fourier analysis of the wave functions follows the three iterations of the Sierpiński triangle. In other words, the self-similar nature of the fractal in the real space is also present in momentum space, see Figs. 6e-h.
Some recent research work has shown an alternative method for realizing Sierpiński gasket based on the deposition of Bi on InSb(111)B [163, 194, 42]. The interest in this platform is related to the semiconducting nature of the InSb(111)B substrate, leading to a lower broadening of the electronic levels in the simulated lattice system [132]. The analysis via STM has revealed the formation of various generations of the Sierpiński gasket [163, 194]. A recent experiment in Canyellas et al. [42] has shown the presence of topological modes in these fractals; theoretical calculations indicate that the nature of these topological modes roots down to the intrinsic spin-orbit coupling induced by the fractal confining potential. The same model shows that these topological modes are suppressed by including a competing spin-orbit coupling of Rashba type [42, 28].
Molecular self-assembly.
Quasicrystals and fractals can also be obtained via molecular self-assembly on a surface. Quasicrystal structures resembling the Penrose tiling have been obtained by depositing C60 molecules on a quasicrystal substrate as Al-Cu-Co [244], Al-Cu-Fe [88] and Bi on Ag-In-Yb [109]. In the case of the fractals, the first notable success is the work of Newkome et al. [188] where three iterations of the Sierpiński hexagonal gasket were realized, and STM identified the final structure. In Shang et al. [236], the Sierpiński gasket has been realized, and the STM has been used to estimate the Hausdorff dimension to a value of .
Twisted 2D layers.
Stacking and twisting layers of 2D materials is a natural way to obtain large-scale quasicrystalline structures. A quasiperiodic lattice with 12-fold rotational symmetry is obtained from twisting two hexagonal lattice structures by an angle of 30∘ relative to each other [249]. This quasicrystal has been realized with twisted graphene bilayers [5, 306], and twisted transition metal dichalcogenides [161]. It has then been proposed that adding a further twisted layer would permit the combination of the flat-band physics at small twist angles, and the quasiperiodic structure appearing at larger angles can be combined [271]. Such quasicrystal flat bands have then been observed in twisted trilayer graphene [107], providing a feasible setup for strongly correlated phases of quasicrystals.
Circuit QED.
Superconducting qubits can implement Bose-Hubbard models with programmable on-site chemical potential and programmable hopping terms. This has been used in Roushan et al. [219], Li et al. [156] to build short chains (up to 10 sites) with quasiperiodic potential and/or tunneling, implementing 1D quasicrystals of bosons.
II.3.2 Lattices with curvature
Lattices with negative curvature have been simulated within circuit QED and classical electric circuits.
Circuit QED.
In Kollár et al. [139], a hyperbolic kagome lattice with Schläfli numbers has been realized in coplanar waveguide resonators coupled to each other via superconducting qubit. Here, the authors have shown that the theoretical transmission spectrum of the circuit is compatible with a -hyperbolic lattice [37], with . This lattice is one of the possible hyperbolic analogs of the kagome lattice and also presents a flat band. For odd, it presents an energy gap between the flat band and the remaining bands absent in the Euclidean case.
Classical electric circuits.
In Lenggenhager et al. [152], the normal modes of Euclidean and hyperbolic lattices have been studied. Specifically, the authors have implemented a lattice with and a lattice with . They have identified a clear difference between the Euclidean and the hyperbolic drum eigenmodes of the two lattices by evaluating the difference in the propagation of waves inside the lattice, see Fig. 7. A similar experimental implementation in Zhang et al. [314] has proposed two different topological implementations of a hyperbolic lattice. In the first one, the authors implemented a hyperbolic version of the Haldane model, showing the existence of chiral edge modes. In the second one, the authors implemented a dimerization of the hopping terms to study higher-order modes; within this approach, they have shown a fractal-like behavior of midgap higher-order zero modes [199]. In a more recent implementation, it has been possible to implement an hyperbolic circuit with periodic boundary conditions and tunable complex phases, as well as hyperbolic graphene lattices showing the Dirac cones [51].
II.3.3 Higher dimensional physics
Classical electric circuits are also a versatile platform for implementing higher-dimensional physics. The two-dimensional array of capacitors and inductors can be easily extended to realize periodic 3D and 4D circuits on a breadboard [315]. The idea is to decompose an D structure into a set of connected D circuit boards; see Fig. 8.
In Wang et al. [288], a 4D lattice has been implemented to realize a topological insulator. Here, the 4D lattice system has been implemented using capacitive and inductive connections fixed in a specific way to implement nonlocal connections in 3D space. Spinless particles with preserved time-reversal symmetry are simulated. Thus, the system is in class AI according to the 10-fold classification of topological metter [223]. Within this classification, a system in class AI presents a topological phase in 4D. Impedance measurements reveal the signature of corner modes associated with this topological phase. In Yu et al. [309], a similar setup is theoretically investigated, and in addition to corner modes, the authors also predict topological signatures as 3D Weyl states.
II.4 Summary
As demonstrated by the above sections, for each of the exotic geometries considered in this colloquium, various possible implementations exist that use atoms, photons, or electrons. Despite the versatility of each platform, they have strengths and weaknesses, which we outline in Tab. 2. Thereby, one should bear in mind that, at least to some extent, these categories are rather subjective, and a nuisance in one context can be an opportunity in another one.
III Applications - novel phenomena in exotic geometries
This colloquium section shall provide an overview of the fascinating topics that can be addressed by implementing the different types of exotic lattices discussed in the previous section. These phenomena interest researchers from various fields, from condensed matter physics to cosmology.
III.1 Localization phenomena
III.1.1 Single-particle localization
Anderson [12] demonstrated that single electrons in uncorrelated static disordered materials exhibit vanishing diffusion at long times, with an exponentially localized wavefunction. The key mechanism in 1D and 2D systems is destructive wave interference everywhere except at the initial point, leading to inevitable localization even in the presence of infinitesimal disorder. However, in 3D, complete backscattering is less likely, requiring a stronger disorder strength for localization to occur, and giving rise to a metal-insulator transition at a critical disorder strength. The transition is characterized by a power-law divergence of the localization length near the critical energy, that is, the mobility edge. In recent times, the phenomenon of localization has also been pursued in the context of quasiperiodic potentials, revealing a rich array of phenomena that challenge our conventional understanding. Compared to disordered systems, quasiperiodic systems display a more diverse localization behavior, including multifractal states and mobility edges even in one dimension. This diversity arises from the unique interplay between disorder and the inherent periodicity of quasiperiodic structures, providing an alternative perspective on the fundamental mechanisms of localization.
Measures of localization.
Let us first introduce some tools to quantify localization properties. A simple and frequently used quantity is the inverse participation ratio of a state
with being the amplitude of the state at site , and being the total number of sites in the system. The mean inverse participation ratio is obtained by averaging over all eigenstates, . For a perfectly localized system, with all states pinned to a single specific site, becomes 1, whereas for perfectly extended (delocalized) states . Hence, the scaling of with is a useful indicator for differentiating localized/extended states. Another classification tool is the -fractal dimension [106, 84] of the eigenstates. For a specific state, this quantity is defined by the following relation . For , the -fractal dimension is proportional to the logarithm of the inverse participation ratio of a state. Averaging over all eigenstates, the mean fractal dimension approaches the value 1 (0) for extended (localized) systems, whereas intermediate values suggest multifractal behavior. The information carried by the -fractal dimensions can also be expressed through the multifractal spectrum , where and [106].
Localization properties of the AA model.
The AA model (see Eqs. (1) and (2)) exhibits all three phases (extended, localized, and critical) [256, 115]. For disorder strength (), all eigenstates are localized (extended). The eigenfunctions are critical or multifractal at .
Without explicit calculations, some of the localization behavior of the AA model can be understood from a duality relation: Consider the Fourier-like transformation which transforms the real-space localized operator into a fully delocalized one, and vice versa. Under this transformation, the AA Hamiltonian (with ) becomes
For , the model is self-dual, that is, the Hamiltonian does not change its form under the duality transformation. At this point, all eigenfunctions must have the same distribution in both real and momentum space.
The localization properties of the AA model have been probed using quasiperiodic photonic waveguide lattices [148] and ultra-cold atoms [216]. The localization transition was identified by analyzing transport, spatial and momentum distributions, and confirmation of scaling behavior for critical disorder strength. These experiments unequivocally demonstrated the transition from extended to localized states as the strength of the quasiperiodic potential increases.
Generalized AA models with mobility edges.
Like 1D disordered systems, the AA model has no mobility edge, and all states in the spectrum behave similarly, i.e., they are either localized or delocalized, irrespective of their energy. However, a 1D quasiperiodic system generally may have mobility edges if the self-duality condition mentioned earlier becomes energy-dependent. In this case, all eigenstates up to a certain energy are localized, whereas states above this “mobility edge” are extended. Different modifications of the AA potential are known to produce a mobility edge [246, 65, 115, 97], for instance:
It contains the AA potential of Eq. (2) as a limiting case () but yields a model with a mobility edge away from this limit. Generalized dual transformations can also be applied to generalized AA (GAA) models with mobility edge [97]. Other ways to obtain a mobility edge in an AA-like model include hopping beyond nearest-neighbors [30, 70]. An additional richness to the localization properties of 1D quasicrystal can be obtained by staggered hoppings, as in the Su-Schrieffer-Heeger model [253]. Increasing the quasiperiodic modulation can immediately delocalize already localized states, giving rise to re-entrant localization transitions [220].
Following Boers et al. [36], Li et al. [158], a 1D quasicrystal with an energy-dependent mobility edge has been implemented with cold atoms in an optical lattice, see Lüschen et al. [170]. An alternative atomic realization of the generalized AA model has been presented in An et al. [8] using a momentum-space lattice. It has also been shown that a balanced bichromatic optical lattice, where the amplitude of both frequency components are comparable, leads to a model with a mobility edge, beyond the standard AA model.
AA model with pairing term.
The presence of pairing terms modifies the localization behavior of the AA model. In this case, the quasiperiodic modulation can also affect the pairing, as studied in Yahyavi et al. [303]. It has been shown that topological superconductivity can co-exist with critical localization properties. While in the absence of pairing the AA model is multifractal only at the critical point of the localization transition, the pairing gives rise to a broader critical regime [89]. The phase diagram becomes even richer in the presence of long-range hopping and/or long-range pairing, where the spectrum exhibits a crossover between localized, multifractal, and ergodic regions [89]. A mixed spectrum has also been seen in the case of the 2D AA-like model, exhibiting extended states on both sides of the self-dual point [254].
Critical behavior of the Fibonacci chain.
In contrast to the AA model, the Fibonacci chain (see Sec. I) does not exhibit a localization transition and remains critical [196, 138] for any value of the modulation strength as dictated by the Luck criterion of bounded fluctuations in the pattern [166]. While the fluctuations can become arbitrarily large for systems with uncorrelated noise, fluctuations are bounded for many quasicrystals, including the Fibonacci quasicrystal. Systems with bounded and unbounded fluctuations have been shown to belong to different universality classes. Using transfer matrix method [25], scaling analysis [166], real-space renormalization group [120, 4], the critical behavior of quantum Ising chains with Fibonacci modulation has been studied, establishing conditions for the occurrence of magnetic phase transitions, and for the quasiperiodic system to be in a different universality class than the uniform one [166, 4]. Also the critical behavior of the Heisenberg model, with quasiperiodic couplings has been studied [113, 114, 121, 4], as well as the XXZ model [277]. It has been shown that a Fibonacci modulation of the Heisenberg coupling alters its critical behavior, in contrast to the XY model or the quantum Ising model, where a quasiperiodic modulation of the coupling is marginal or irrelevant as the central charge remains unaffected. However, we come to these later since these spin models map to interacting fermionic models.
Localization in 2D quasicrystals.
Non-interacting systems in lower-dimensional landscapes, defined by static, random potential energy and short-ranged tunneling, exhibit Anderson localization. Transitioning to higher dimensions facilitates delocalization, driven by increased energy density. For quasicrystals, with the increasing sophistication of experimental techniques, research into localization has extended to two-dimensional (2D) quasicrystals, revealing a multifaceted landscape of localization behavior. Quasicrystalline order usually arises from projecting a higher-dimensional periodic lattice in an incommensurate manner. Penrose introduced a method for constructing quasicrystals using a set of tiles and matching rules, resulting in the well-known fivefold symmetric Penrose tiling and the eightfold symmetric octagonal tiling [202]. These tilings, with disallowed rotational symmetries, exhibit self-similarity in both real and reciprocal space. Self-similarity upon scaling indicates nontrivial structure at arbitrarily large scales, reflected in quasicrystal diffraction patterns with sharp peaks at very small momenta. In the studies Viebahn et al. [276], Sbroscia et al. [228], a BEC of Potassium atoms is utilized to probe a 2D quasicrystalline optical lattice through stimulated two-photon Kapitza-Dirac scattering. The quasicrystalline potential is created by combining four mutually incoherent 1D optical lattices, forming a global eightfold rotational symmetry. Unlike periodic lattices, the combination of reciprocal lattice vectors results in new, smaller momentum scales, creating a self-similar fractal structure (see Fig. 1b). This process is akin to the incommensurate projection of a 4D simple cubic lattice to the 2D plane, similar to the construction of the octagonal tiling. Adjusting the number of lattice beams allows control over the dimensionality of the parent lattice. It has been predicted that the spectrum consists of a localized low-energy part, an extended high-energy part, and an intermediate regime, in which localized and critical states alternate [318]. Finally, adiabatically loading BECs into quasicrystal lattices allows studying Bose glasses in this geometries [308]. The system thermalizes at higher energy density and weaker disorder, exhibiting a conducting phase.
Localization in fractal lattices.
Also, fractal structures, such as the Sierpiński gasket, are known to host localized eigenstates [73], in addition to an infinite number of extended states [286]. This can give rise to complex transport behavior, featuring an increased return probability in quantum walks as compared to classical random walks [63]. Recently, quantum transport in fractal geometries has been explored in waveguide experiments [301], and super-diffusive quantum transport has been observed in both Sierpiński gasket and carpet. However, in the experimental setup connectivity between all nearest neighbors has avoided a bottleneck in the Sierpiński gasket, which occurs at sites where two fractal generations are glued together. As pointed out by Ref. Rojo-Francàs et al. [218], this bottleneck significantly affects the energy spectrum of the system, and slows down quantum transport into the sub-diffusive regime. Theoretical studies of fractal lattices with random connectivity have shown that eigenfunction localization depends on the spectral dimension rather than the Hausdorff dimension [143].
Localization in higher-dimensional or curved spaces.
While quasicrystalline or fractal geometries may give rise to localization in the absence of disorder, quantum particles in Euclidean 1D or 2D lattices are localized by an infinitesimal amount of disorder. A different behavior has been reported for hyperbolic lattices, where finite disorder is required for localization [52], similar to the localization behavior in higher dimensions , see [261]. Quantum funneling, a localization effect caused by the singularity in a negatively curved space, has been described by Zhang et al. [312]. In the future, quantum simulators of exotic geometries may be used to probe localization behavior in higher-dimensional or curved spaces.
III.1.2 Many-body localization
Quantum simulators of quasicrystals have also yielded groundbreaking discoveries in the many-body context, showcasing their ability to test theoretical predictions experimentally. Traditionally, a disordered potential localizes single-particle quantum eigenstates, rendering the system an insulator with zero conductivity. In Basko et al. [22], it was proposed that interacting many-body systems could similarly undergo a MBL transition. This transition involves highly excited states, occurring at low energy density and strong disorder, where the system remains a perfect insulator. The system thermalizes at higher energy density and weaker disorder, exhibiting a conducting phase. MBL violates the eigenstate thermalization hypothesis (ETH), which posits that each eigenstate at finite energy density has the same expectation values of local physical observables as the ones provided by local thermal states. This implies that thermalization occurs at the level of individual eigenstates in an isolated quantum system, where no memory of the initial state survives. The eigenstates of the system in the MBL phase do not obey the ETH, and some memory of the initial conditions is retained in local observables for arbitrarily long times.
Theoretical investigations.
Various theoretical studies have considered localization on the many-body level in quasiperiodic potentials. In fact, the quasiperiodic scenario with AA modulation and the purely random scenario have explicitly been compared in Khemani et al. [133], and it has been argued that the quasiperiodic potential produces a more stable MBL phase than the quenched disorder. Moreover, it has been claimed that the two systems belong to different universality classes. A huge number of theoretical works addressed the MBL phase of interacting particles in the AA potential [124, 133, 26, 153, 235, 151, 292, 72, 300, 305]. There is full agreement that, in the case of an AA modulation, the localization transition in interacting systems occurs at larger values of the quasicrystal potential than in the free system. Most of the literature results locate the transition within . However, no narrow value exists. One reason for this is that all studies are finite-size studies, which leads to some intermediate regime, cf. Rispoli et al. [215] and Fig. 9(d). However, in this context, we also note that Doggen and Mirlin [72] claims that finite-size effects are rather weak for MBL systems in quasiperiodic potentials, which is in contrast to the purely random disorder. On the other hand, Xu et al. [300] studies the scrambling of a localized perturbation and finds that it is suggestive of an intermediate phase. Another reason why the transition value is not sharply determined is that different figures of merit can be used to determine the localization transition. For instance, Lee et al. [151] reports by considering the finite-size scaling of the entanglement entropy, but , from the quench dynamics, when the system is initially prepared in a very imbalanced state. There have also been theoretical studies of localization behavior in interacting Fibonacci chains [172, 274, 55], as well as in the interpolating AA-Fibonacci potential [282]. While a non-interacting Fibonacci system remains critical for any potential strength, that is, it does not localize, Macé et al. [172], Chiaracane et al. [55] claim evidence of a crossover to MBL for Fibonacci chains with sufficiently strong interactions. On the other hand, the absence of a MBL phase was reported for a weakly interacting Fibonacci system [274]. Recent theoretical work has also studied MBL on fractal lattices in the presence of random disorder [178]. These systems provide an excellent opportunity for exploring the still unknown fate of MBL beyond 1D systems.
In general, the theoretical description of the MBL transition is extremely challenging. In contrast to the metal-insulator transitions, which occur in the ground state or the low-energy states of a system, the MBL transition affects the whole spectrum or at least an extended spectral range, which might be far away from the ground state. For this reason, much of the machinery to study many-body phases of matter (such as Monte Carlo techniques or tensor network methods) is not well suited for studying MBL. As a consequence, numerical explorations have mostly relied on exact diagonalization studies of systems with 20-30 spins, and even these relatively moderately sized studies strongly profit from sophisticated techniques [204, 241, 240], exploiting the sparseness of the matrix. Larger system sizes have been treated with a self-consistent Hartree-Fock description [292] or time-dependent density matrix renormalization group (DMRG) studies [153, 300].
Different fingerprints of the MBL phase are known: (i) The half-chain von Neumann entanglement entropy is strongly suppressed in the MBL phase [151, 153, 172]. The transition point can be estimated by considering the value at which the entanglement entropy per site is independent of the system size . (ii) The level spacing ratio , where is the spacing between adjacent energy levels, differentiates between localized and ergodic phase. It is well established by random matrix theory that the level spacing distribution characterizes the statistical ensemble to which a matrix belongs [105]. Ergodic systems typically belong to the Wigner-Dyson ensemble, characterized by level repulsion and a relatively large value . On the other hand, integrable/localized systems are characterized by the Poisson statistics, , see Pal and Huse [200]. (iii) Participation ratio and fractal dimension of the eigenstates, as introduced in the Section on non-interacting systems, can also be used for a many-body system if, instead of the full many-body eigenstates, the eigenstates of the one-body density matrix are considered [26, 172]. (iv) Dynamics probes have been used in many experiments: If a system is localized, it should keep memory of its initial state during a quench. Often, systems are prepared in a state with alternating empty and occupied sites, and the evolution of this even/odd population imbalance is tracked. In contrast to an ergodic system, the MBL system remains imbalanced on long time scale [232].
Experimental studies.
Given the computational difficulties, quantum simulation has become an extremely valuable tool for studying MBL. At the same time, due to the instability of the MBL phase against a thermal bath, the study of the MBL phase is also motivated by the unique opportunity of isolating quantum systems from the environment offered by quantum simulators. The primary systems where the MBL transition is experimentally studied are cold atoms in optical lattices with an AA modulation [232, 169, 137, 167, 215]. Experiments in Munich [232, 169, 137] have concentrated on rather larger systems, arrays of 1D tubes with approximately 150-200 sites per tube, filled with spin-1/2 fermions.
An important figure of merit in these experiments is the population imbalance between even and odd sites. Initially, the system was prepared for a large imbalance, which was then washed out through dynamic evolution. While in the absence of a quasiperiodic potential, the imbalance fully disappears on short time scales, a sufficiently strong quasiperiodic potential leads to saturation of the imbalance at a finite value, see Fig. 9a. By plotting the saturation value versus the potential strength , see Fig. 9b, the onset of localization is found to be shifted towards larger values of by the interactions. In the non-interacting case, the transition seems to occur slightly below the single-body critical value , which can be accounted for by the harmonic trapping potential compared with exact numerical simulations. In the vicinity of the localization transition, especially in the interacting system, the saturation occurs only on much longer time scales than deeply in the ergodic or the MBL regime. Kohlert et al. [137] also studied MBL in a GAA model. In contrast to the AA model, the non-interacting GAA model exhibits an intermediate phase in which both localized and delocalized eigenstates co-exist. While the experiment confirmed the MBL behavior of the interacting GAA, a many-body counterpart of the intermediate phase has not been seen.
The experiments in Harvard, in particular Rispoli et al. [215], have concentrated on this critical regime, measuring observables like correlations between sites, entropy, and particle number fluctuations. The experimental procedure of Lukin et al. [167], Rispoli et al. [215] starts from a bosonic Mott insulator with one atom per site and measures the particle number per site after some evolution time. From this, different quantities can be computed. For instance, the probability of finding particles on a site determines the single-site von Neumann entropy, . For weak quasiperiodic potentials, the obtained value matches well with the expected thermal average, but it becomes reduced when the quasiperiodic modulation becomes stronger, approximately for , see Fig. 9c, notably beyond the critical value of the single-particle AA model. In contrast, the entropy reaches a maximum within the transition regime as the system strongly fluctuates. Criticality is also characterized by strong -site correlations , shown in Fig. 9(d).
Bosonic systems have also been produced with up to 10 superconducting qubits, realizing Bose-Hubbard chains with an AA-like potential [219], and/or AA-modulated hopping [156]. All energy levels can be detected if prepared with only one or two bosons, revealing the fractal butterfly spectrum in the single-particle case. Using superconducting qubits, it has also been demonstrated that the level spacing of an interacting system is shifted towards smaller values by increasing the quasiperiodic modulation strength. By addressing bosonic systems, quantum simulators may finally help to solve the question of whether MBL of bosons exists [58].
III.2 Topological phenomena
Topological quantum systems are characterized by a band structure with non-trivial topological properties, as in the case of topological insulators [211], or by a topological many-body ground state, as in the case of FQH systems and topological superconductors [187]. In all cases, the non-trivial topological properties are mathematically defined via non-local quantities, specifically non-zero Chern numbers, which are physically linked to robust properties, such as transport behavior or fractional quantum statistics of quasiparticles. Topological matter has been an intense research field in the past decades, mainly in the context of systems in “regular” geometries and driven by the search for robust quantum systems. Quasicrystalline or fractal structures and higher-dimensional lattices provide interesting new aspects to the field of topological matter, as discussed in the following.
III.2.1 Charge-pumping, Chern numbers, and gap labelling in 1D quasicrystals
There is a remarkable formal equivalence between AA-type 1D quasicrystals and the Harper-Hofstadter model [108, 116]. The latter describes electrons in a 2D periodic lattice subject to a uniform magnetic field along the axis. In the Landau gauge, , the Harper-Hofstadter Hamiltonian is diagonalized along the -direction via a Fourier transform, resulting in an effective AA potential along the -direction, with the phase parameter given by the wavevector and the modulation period by the magnetic flux. Hence, the magnetic field leads to a quasiperiodic structure when the lattice constant and magnetic length are incommensurate. This formal equivalence permits the use of topological concepts established for 2D crystals to be applied to 1D quasicrystals. In particular, as a paradigmatic model for the integer quantum Hall effect, the Harper-Hofstadter model has played a key role in connecting topological Chern numbers of the energy bands in the periodic system to quantized transport phenomena [263, 95]. These integer numbers are obtained by considering a gapped band, that is, a set of states which are energetically separated from other states and which extend over the whole 2D Brillouin zone of the system, and integrating over the Berry curvature of this set of states:
Here, repeated indices are summed, is the antisymmetric tensor, , and the Berry connection . Physically, the Chern number quantifies the transverse conductivity of the band [263], as well as the number of states which, in the case of open boundary conditions, are localized at the system edge and connect different bands by winding through the energy gap [110]. This relation between bulk transport properties and edge state is known as bulk-edge correspondence. With the Chern number being the band’s topological property, the related physical behavior is robust against local perturbations.
In the context of an AA model, Chern numbers are obtained by substituting with the phase angle in the quasiperiodic potential. Such a procedure can be applied to arbitrary 1D models with a cyclic parameter [144, 146, 275, 104]. Chern numbers of 1D quasicrystals have established a topological equivalence between different models, specifically between the AA model and a Fibonacci chain [146, 275]. On the other hand, in the same way as tuning a magnetic field in a 2D lattice produces changes to the Hall conductivity, tuning the quasiperiodic modulation parameter in the AA model produces topological phase transitions in the 1D quasicrystal [275]. The robustness of the Chern numbers against on-site interactions has been studied in Matsuda et al. [180].
In the case of 1D quasicrystals, the physical relevance of non-zero Chern numbers is evidenced by pumping experiments, where a quantized amount of “charge” is pumped through the chain when the cyclic Hamiltonian parameter is adiabatically changed. The amount of the pumped “charge” is defined by the Chern numbers of the filled bands. Such topological charge pumping has already been proposed by Thouless [262] in the context of 1D periodic (super)lattices. It has been realized in different photonic quasicrystals built from coupled waveguides in Kraus et al. [144], Verbin et al. [275] by adiabatically sweeping the phase of the quasiperiodic potential along the propagation axis. This setup has also served to visualize the presence or absence of localized states at the boundary between different quasicrystals [275], depending on whether the quasicrystals are topologically equivalent or not, see Fig. 10a. With cold atoms, one can fill the lowest band either as a fermionic band insulator [185] or a bosonic Mott insulator [164] and obtain a quantized charge transport of one site per modulation cycle.
From a static point of view, that is, without change of a Hamiltonian parameter, the topology of 1D quasicrystals is captured by the gap labeling theorem [24, 23]. This theorem generally applies to 1D Schrödinger equation with bounded potential, stating that within the energy gaps, the integrated density of states takes quantized values expressed by integer labels. For instance, for a Fibonacci chain, is given by
where is the golden mean, and are integers. Like the Chern numbers, these gap labels and are topologically robust: a perturbation does not change them unless the perturbation closes the band gap. The accuracy of the gap labeling theorem in Fibonacci chains has been demonstrated experimentally using cavity-polaritons [255, 15]. These observations suggest that Chern numbers and gap labels are two sides of the same coin [64].
III.2.2 4D quantum Hall effect in 2D quasicrystals
The mapping between 1D superlattices and 2D quantum Hall models can be extended to higher dimensions: Modulating a 2D lattice along both directions, it can be mapped onto a 4D quantum Hall system [145]. In such a system, a Hall current is induced by the combination of a perpendicular electric field perturbation and a magnetic field perturbation in the -plane perpendicular to and . Such a response is described by novel topological concepts, specifically the second Chern number. It can be obtained by decomposing the (2+2)-dimensional system into two (1+1)-dimensional systems along each physical dimension . The Berry curvature of the (1+1)-dimensional system along is defined through the wavevector and the modulation . The second Chern number is then obtained by integrating the product of Berry curvatures over the 4D Brillouin zone [210, 182],
If the bands are factorizable along and , the second Chern number can be decomposed into products of the first Chern numbers. This decomposition then also applies to the edge states: These states can either be 0D corner states, obtained as a product of the 0D edge states of the individual 1D systems, or a 1D state, obtained as the product of a 0D edge state and a 1D bulk state.
Experimentally, these different aspects of the 4D quantum Hall effect have been probed with atomic and photonic quantum simulators [165, 319]. Specifically, 2D photonic waveguides with modulated coupling along the propagation direction have probed 4D quantum Hall response via photon pumping from edge to edge and corner to corner [319], see also Fig. 4. Cold atoms in an optical 2D superlattice have been used to study the bulk behavior of a 4D quantum Hall system [165]. This setup has allowed for measuring the second Chern number via the response of the atoms to a perturbing electric field, mimicked by a phase shift of the superlattice potential, and to a magnetic perturbation, mimicked by a small tilt of the long lattice with respect to the corresponding short lattice.
We also note the recent implementation of the 4D quantum Hall effect in a system with synthetic extra-dimensions encoded in the large electronic spin manifold of dysprosium atoms [39].
III.2.3 Topological classification
The ten-fold topological classification [223] considers the behavior of a Hamiltonian under three fundamental discrete symmetries: time-reversal , charge-conjugation , and the product of the two . The ten-fold way establishes that a system that breaks all these symmetries can have non-zero Chern numbers only in even dimensions. The purported contradiction of this scheme through 1D quasicrystals is resolved by noting that the Chern numbers for the quasicrystals do not classify a single 1D Hamiltonian, but a parametrized family of models [173]. From this perspective, the ten-fold topological classification remains unchanged for quasicrystals.
Beyond the ten-fold way, there are also crystal symmetries that can protect topological phases. This can give rise to weak topological insulators/superconductors or higher-order topological insulators (HOTIs) [57]. In total, there are 528 (1651) magnetic space group symmetries in 2D (3D), which can be used for classifying topological quantum matter by crystal symmetries [290]. However, since quasicrystals, by definition, support symmetries that cannot exist in crystals, it is not surprising that topological phases without crystalline counterparts become possible in quasiperiodic systems [273, 53, 82, 248]. In Varjas et al. [273], it has been shown that a HOTI with gapless topological corner modes exists for a 2D quasicrystal with eightfold rotational symmetry. A classification scheme for topological phases in quasicrystals has been presented in Else et al. [82]. Still, the topological richness of quasicrystals has not yet been fully explored. HOTIs can also occur in higher-dimensional spaces. In Dutt et al. [74], a 4D HOTI has been realized using the synthetic frequency dimension of photonic molecules.
III.2.4 Topological edge states in fractals and quasicrystals
New topological phases are also expected or already observed in fractal and hyperbolic lattices. Whereas for standard geometries, the bulk-edge correspondence establishes an important relation between topological bulk behavior (Chern numbers) and topological edge states, extended bulk regions are lacking in fractal systems. Non-interacting topological systems (i.e., topological insulators and superconductors) on Sierpiński fractals have been studied theoretically in Brzezińska et al. [189], Pai and Prem [199], Iliasov et al. [122], Fremling et al. [94] , and in the Penrose tiling [86]. In these systems, the Hall conductivity is neither quantized nor proportional to the Chern number of the filled bands. However, splitting into various sub-bands makes it difficult to discern whether a band is filled. The fate of edge states is particularly intriguing since, on a fractal lattice, most sites are near some edge of the system. Topological edge states in a Sierpiński gasket have recently been observed in a photonic setup [31, 157], see Fig. 10(b).
Anyons are the hallmark of interacting topological systems. The question of how these exotic quasiparticles are affected by fractal dimensions has been theoretically addressed in Manna et al. [179, 177], giving exact parent Hamiltonians for analytic FQH wave functions on Sierpiński fractals and demonstrating anyonic braiding behavior of their quasiparticles. The entanglement entropy scaling in fractional Chern states on fractal lattices has been studied in Li et al. [160], and a violation of the area-law indicates that entanglement entropy is not suited to characterize topological phases on fractal lattices.
Theoretically simpler than the FQH effect are topological superconductors, such as the paradigmatic Kitaev chain [135]. Even on a mean-field level, these systems support non-Abelian anyons, specifically Majorana zero modes localized at the ends of the chain. These quasiparticles, which are their own antiparticles, possess intriguing topological properties that can be used to encode and process quantum information. Signature of Majorana modes have been observed in chains of magnetic atoms placed on top of superconductors [184, 231, 307]; however, there is no consensus on the topological origin of the observed boundary modes [147]. Following the work by Soldini et al. [245], where 2D regular lattices of magnetic impurities were created on a superconductor, we envision that this technique could be extended to construct quasicrystal and fractal lattices of magnetic impurities to realize topological superconductors in non-standard geometries. Quasiperiodic models, and especially the AA model, have attracted a lot of attention in the context of topological superconductivity [41, 68, 69, 311, 284, 303, 90]. The quasiperiodic potential provides an excellent testing ground for the robustness of the topological phase, as it does not break the protecting symmetries. Both analytic and numerical studies have explicitly confirmed that the topological phase and the Majorana quasiparticle persist in the presence of a weak quasiperiodic potential, but a strong incommensurate modulation drives a localization transition that destroys the topological phase. Notably, in some cases, quasiperiodic modulation of the chemical potential [90] or of the hopping [284] might even stabilize the topological phase.
The recent realization of hyperbolic lattices has inspired the development of hyperbolic lattice crystallography [38], including topological phases such as hyperbolic Haldane and Kane-Mele models [272, 314]. A relevant feature for topological phases is the extensive scaling of the boundary for hyperbolic lattices, where a macroscopic fraction of all states contributes to topological edge states. Hyperbolic lattices can feature non-trivial second Chern numbers for zero first Chern numbers [313] as well as higher-order topology with an unconventional number of corner modes, not allowed in crystalline materials [260]. A hyperbolic photonic topological insulator was recently realized using coupled ring resonators on silicon chips [118].
III.3 Cosmological models
The engineering of systems with curved geometries provides a new tool for studying gravitational physics and cosmology in the laboratory. As examples, we first consider the quantum simulation of gravitational effects in lattices, which explicitly implement curvature through the appropriate connectivity design. We then also mention the wide field of analog gravity: The analogy between sound propagation on a background hydrodynamic flow and field propagation in curved spacetime, first pointed out by Unruh [270], allows for studying gravitational phenomena even in flat systems. A similar analogy can also be made between light propagation in inhomogeneous media and field propagation in curved spacetime [134].
The Unruh effect [96, 66, 269] and the Hawking radiation [111] are two important aspects related to the physics of black holes. Both phenomena are based on predictions of quantum field theory in curved spacetimes and non-inertial reference frames. The Unruh effect predicts that an observer with constant acceleration through empty space will perceive a thermal bath characterized by the Unruh temperature , where is the speed of light in vacuum, is the reduced Planck constant, and is the Boltzmann constant. The same is not true for an inertial observer. Hawking predicted that black holes are far from being emission-free, and they should emit a steady flux of thermal radiation, known as Hawking radiation, with a temperature proportional to , the gravitational field strength at the event horizon . Both effects are extremely weak and have not been observed using astronomical techniques. In this respect, analog quantum simulators of these are very appealing.
For the case of the Unruh effect, Rodríguez-Laguna et al. [217], Kosior et al. [142] proposed an implementation with cold atoms in an optical lattice. The general idea is to simulate a tuneable Dirac-Hamiltonian system that can interpolate between a Minkowski (rest frame) and a Rindler (constant accelerated frame) configuration. Specifically, the key difference between the two implementations is the gauge field characterizing the cold atoms’ hopping terms in the optical lattice. These two different Hamiltonians are implemented, e.g., in a square lattice with constant hopping term along the two directions and by adding a gradient diagonally to the lattice. The difference in the hopping terms between the Minkowski and the Rindler implementation can be obtained by applying an additional Raman laser collinear with the lattice beams. Finally, information about the Unruh effect is obtained by measuring the Wightman response, which is a two-point correlation function [217]. A similar implementation for interacting cold atoms is proposed in Kosior et al. [142].
There is also an analogy between the general relativity and the Hamiltonian of Weyl’s systems [280]. A Weyl Hamiltonian can be obtained as low-energy excitations of Weyl semimetals [14]. The Hamiltonian describing Weyl semimetals reads
where is the Fermi velocity, the three-dimensional momentum, and the vector of Pauli matrices. For small tilts, , the Weyl cone is denoted as type I. For large tilts , the Weyl cone is overtilted and called type II [14]. Type I Weyl cones correspond to light cones in a flat Minkowski spacetime, whereas type II Weyl nodes to light cones that tilt close to black or white holes. The former is for tilting towards the horizon, and the latter is for tilting away from the horizon — see Fig. 11(a-b). There is a simple mapping between Eq. (12) and the spacetime metrics: one must define the frame fields and their inverse [280, 281]. In general relativity, a frame field defines a local orthonormal coordinate system at each point in space-time. They are connected to the spacetime metric as , where represents the Minkowski metric. The terms in Eq. (12) can be rearranged to introduce frame fields as the tensor connecting Pauli matrices and momenta; in this way, we obtain:
with . In Eq. (13), Roman (Greek) indices run over spatial (space-time) coordinates, and we are assuming the summation over identical indices. The essence of the mapping between Weyl Hamiltonians and black holes is that the frame fields defined from Eq. (13) coincide with the ones of the Gullstrand-Painlevé-Schwarzschild metric associated with , with the tilt taking the role of the velocity . Thus, a Weyl Hamiltonian with a continuous tilt changing from to can be mapped to the metric of a black or white hole in Gullstrand-Painlevé coordinates. The location of the critical tilt, , is the analog of the event horizon.
In De Beule et al. [67], Sabsovich et al. [224], a continuous smooth junction between a type I and type II Weyl semimetal has been proposed to study an artificial event horizon. Transport experiments can give rise to the signature of Hawking-like radiation [67]. When considering a more realistic implementation of Weyl semimetals in a lattice system, there is a doubling of the Weyl cone. Consequently, this gives rise to new effects dubbed Hawking fragmentation and Hawking attenuation of a wave packet approaching the event horizon [224]. Junctions of this type cannot be implemented in realistic materials hosting Weyl semimetals but could be engineered in acoustic crystals [304, 203], photonics [119], topoelectrical circuits [212], and cold atoms [302]. A recent proposal by Stålhammar et al. [252] extends these results to the realm on non-Hermitian symmetric systems.
The exceptional tunability of certain quantum systems, notably cold atomic systems and quantum fluid of light (i.e., non-linear transport of polariton BEC), renders them ideal for simulating particle flows that mimic curved spacetime for the system’s phononic excitations. For instance, the concept of “sonic black holes” within a BEC was proposed [98, 16, 48], leading to the realization of a polaritonic black hole [192] and the subsequent atomic realization of sonic black holes as experimental tools for investigating Hawking radiation [250, 193, 140]. The main idea behind these realizations is similar to the Unruh effect in classical fluids but now applied to quantum fluids: the “sonic horizon” separates a subsonic upstream region () from a supersonic downstream region (), as illustrated in Figs. 11(c,d), where is the speed of the BEC flow and are the speed of sound in the upstream and downstream regions, respectively. To probe the Hawking radiation in the case of 1D black holes, in Balbinot et al. [16], Carusotto et al. [49] has been proposed to use the second-order correlation function of density fluctuation, given by:
with denoting normal ordering. The signature of Hawking radiations, i.e., correlated phonon pairs that are continuously created at the horizon, is a symmetric pair of negative correlation stripes in the mapping with a slope of . This corresponds to the propagation of correlated phonon pairs in the upstream and downstream region at speeds and , respectively. This correlation has been numerically demonstrated using Monte Carlo simulations [49, 192] and experimentally observed in atomic BEC system [250], as shown in Fig. 11(e). Moreover, the ability to modify the curved metric for phonons in a BEC dynamically has facilitated the exploration of cosmological models, providing insights into the equations describing the universe’s expansion dynamics [278].
IV Conclusion and Outlook
This colloquium has aimed at reviewing the state-of-the-art quantum engineering techniques and their reachable physical phenomena from a context that concentrates on exotic geometries. To this aim, we have covered various systems (atomic, electronic, and photonic ones), highlighting unique opportunities offered by different platforms and introducing general concepts such as Floquet engineering and synthetic dimensions. Many of these developments involve quantum control on the microscopic level and, therefore, resonate with strong efforts of building programmable quantum simulators. Such devices may then serve computational purposes and eventually demonstrate a quantum computational advantage, e.g. via boson sampling using a network of photons [316], or via quantum optimization using an Rydberg atoms in optical tweezer arrays [76].
In this colloquium, we have concentrated on specific physical implications of the exotic geometries. This includes (i) localization phenomena, (ii) topological phenomena, (iii) analog black holes, and cosmology. These aspects already contain a rich variety of physical behavior on the single-body level, such as mobility edges in low dimensions (smaller than 3) or integer quantum Hall effect in high dimensions (larger than 3). While some aspects, such as topological classification in non-periodic crystals, remain an outstanding challenge, many features related to single-particle physics in exotic lattices have been theoretically understood and experimentally realized. Much less explored, though, is the complicated subject of quantum many-body physics in such geometries. In this regard, the present colloquium article has provided a brief perspective on many-body localization. Theoretical research on this topic has indeed been ignited by the possibility of experimentally building synthetic quasicrystals that show this intriguing behavior or similar behavior (such as glassy phases). Yet there is open debate on the true existence of MBL phase, particularly in more than one dimension. Very little is known regarding topological many-body phenomena in quasicrystals or fractals. Still, the rich single-particle physics triggers hope for an even richer many-body scenario. Given the complexities in simulating such systems classically, the quantum simulators may develop their full potential of opening the door to deep questions of quantum many-body physics. With the engineering of exotic kinetic terms being achieved now, the basis has been set for future explorations of the interacting world of fractals, quasicrystals, and curved and higher-dimensional spaces.
While this review has focused on quantum simulators where the setup imposes exotic geometries, there are also fascinating opportunities to observe similar geometric structures emerging spontaneously, for instance, due to multi-body interactions. An example are fracton phases, where fractal structures arise as the low-dimensional subspace of exotic quasi-particles with restricted mobility [186, 207]. While the requirement of multispin interactions is difficult to engineer, Rydberg atoms have recently been proposed as a suitable experimental platform [183]. This might open another intriguing avenue towards quantum systems with exotic geometries [279].
Acknowledgements.
[…]
- Abbott [1884] Abbott, Edwin A (1884), Flatland: Romance of Many Dimensions (Seeley).
- Adiyatullin et al. [2023] Adiyatullin, Albert F, Lavi K. Upreti, Corentin Lechevalier, Clement Evain, Francois Copie, Pierre Suret, Stephane Randoux, Pierre Delplace, and Alberto Amo (2023), “Topological properties of floquet winding bands in a photonic lattice,” Phys. Rev. Lett. 130, 056901.
- Agrawal et al. [2023] Agrawal, Shraddha, Sai Naga Manoj Paladugu, and Bryce Gadway (2023), “Two dimensional momentum state lattices,” arXiv:2305.17987 [cond-mat.quant-gas] .
- Agrawal et al. [2020] Agrawal, Utkarsh, Sarang Gopalakrishnan, and Romain Vasseur (2020), “Universality and quantum criticality in quasiperiodic spin chains,” Nat. Commun. 11 (1), 2225.
- Ahn et al. [2018] Ahn, Sung Joon, Pilkyung Moon, Tae-Hoon Kim, Hyun-Woo Kim, Ha-Chul Shin, Eun Hye Kim, Hyun Woo Cha, Se-Jong Kahng, Philip Kim, Mikito Koshino, Young-Woo Son, Cheol-Woong Yang, and Joung Real Ahn (2018), “Dirac electrons in a dodecagonal graphene quasicrystal,” Science 361 (6404), 782–786.
- Albuquerque and Cottam [1992] Albuquerque, EL, and M.G. Cottam (1992), “Theory of plasmon-polaritons in fibonacci-type superlattices with two-dimensionl electron gas layers,” Solid State Commun. 81 (5), 383–386.
- Altman et al. [2021] Altman, Ehud, Kenneth R. Brown, Giuseppe Carleo, Lincoln D. Carr, Eugene Demler, Cheng Chin, Brian DeMarco, Sophia E. Economou, Mark A. Eriksson, Kai-Mei C. Fu, Markus Greiner, Kaden R.A. Hazzard, Randall G. Hulet, Alicia J. Kollár, Benjamin L. Lev, Mikhail D. Lukin, Ruichao Ma, Xiao Mi, Shashank Misra, Christopher Monroe, Kater Murch, Zaira Nazario, Kang-Kuen Ni, Andrew C. Potter, Pedram Roushan, Mark Saffman, Monika Schleier-Smith, Irfan Siddiqi, Raymond Simmonds, Meenakshi Singh, I.B. Spielman, Kristan Temme, David S. Weiss, Jelena Vučković, Vladan Vuletić, Jun Ye, and Martin Zwierlein (2021), “Quantum simulators: Architectures and opportunities,” PRX Quantum 2, 017003.
- An et al. [2021] An, Fangzhao Alex, Karmela Padavić, Eric J. Meier, Suraj Hegde, Sriram Ganeshan, J. H. Pixley, Smitha Vishveshwara, and Bryce Gadway (2021), “Interactions and Mobility Edges: Observing the Generalized Aubry-André Model,” Phys. Rev. Lett. 126, 040603.
- An et al. [2018a] An, Fangzhao Alex, Eric J. Meier, Jackson Ang’ong’a, and Bryce Gadway (2018a), “Correlated dynamics in a synthetic lattice of momentum states,” Phys. Rev. Lett. 120, 040407.
- An et al. [2017] An, Fangzhao Alex, Eric J. Meier, and Bryce Gadway (2017), “Diffusive and arrested transport of atoms under tailored disorder,” Nat. Commun. 8 (1), 1–6.
- An et al. [2018b] An, Fangzhao Alex, Eric J. Meier, and Bryce Gadway (2018b), “Engineering a flux-dependent mobility edge in disordered zigzag chains,” Phys. Rev. X 8, 031045.
- Anderson [1958] Anderson, P W (1958), “Absence of diffusion in certain random lattices,” Phys. Rev. 109, 1492–1505.
- Anisimovas et al. [2016] Anisimovas, E, M. Račiūnas, C. Sträter, A. Eckardt, I. B. Spielman, and G. Juzeliūnas (2016), “Semisynthetic zigzag optical lattice for ultracold bosons,” Phys. Rev. A 94, 063632.
- Armitage et al. [2018] Armitage, N P, E. J. Mele, and Ashvin Vishwanath (2018), “Weyl and Dirac semimetals in three-dimensional solids,” Rev. Mod. Phys. 90, 015001.
- Baboux et al. [2017] Baboux, Florent, Eli Levy, Aristide Lemaître, Carmen Gómez, Elisabeth Galopin, Luc Le Gratiet, Isabelle Sagnes, Alberto Amo, Jacqueline Bloch, and Eric Akkermans (2017), “Measuring topological invariants from generalized edge states in polaritonic quasicrystals,” Phys. Rev. B 95, 161114.
- Balbinot et al. [2008] Balbinot, Roberto, Alessandro Fabbri, Serena Fagnocchi, Alessio Recati, and Iacopo Carusotto (2008), “Nonlocal density correlations as a signature of Hawking radiation from acoustic black holes,” Phys. Rev. A 78, 021603.
- Balčytis et al. [2022] Balčytis, Armandas, Tomoki Ozawa, Yasutomo Ota, Satoshi Iwamoto, Jun Maeda, and Toshihiko Baba (2022), “Synthetic dimension band structures on a Si CMOS photonic platform,” Sci. Adv. 8 (4), 10.1126/sciadv.abk0468.
- Banik et al. [2022] Banik, S, M. Gutierrez Galan, H. Sosa-Martinez, M. J. Anderson, S. Eckel, I. B. Spielman, and G. K. Campbell (2022), “Accurate Determination of Hubble Attenuation and Amplification in Expanding and Contracting Cold-Atom Universes,” Phys. Rev. Lett. 128, 090401.
- Barczyk et al. [2024] Barczyk, René, L. Kuipers, and Ewold Verhagen (2024), “Observation of Landau levels and chiral edge states in photonic crystals through pseudomagnetic fields induced by synthetic strain,” Nature Photonics 18 (6), 574–579.
- Barredo et al. [2016] Barredo, Daniel, Sylvain de Léséleuc, Vincent Lienhard, Thierry Lahaye, and Antoine Browaeys (2016), “An atom-by-atom assembler of defect-free arbitrary two-dimensional atomic arrays,” Science 354 (6315), 1021–1023.
- Barsukova et al. [2024] Barsukova, Maria, Fabien Grisé, Zeyu Zhang, Sachin Vaidya, Jonathan Guglielmon, Michael I. Weinstein, Li He, Bo Zhen, Randall McEntaffer, and Mikael C. Rechtsman (2024), “Direct observation of Landau levels in silicon photonic crystals,” Nature Photonics 18 (6), 580–585.
- Basko et al. [2006] Basko, D M, I L Aleiner, and B L Altshuler (2006), “Metal–insulator transition in a weakly interacting many-electron system with localized single-particle states,” Ann. Phys. (N. Y.) 321 (5), 1126–1205.
- Bellisard et al. [1992] Bellisard, Jean, Anton Bovier, and Jean-Michel Ghez (1992), “Gap labelling theorems for one dimensional discrete Schrödinger operators,” Rev. Math. Phys. 04 (01), 1–37.
- Bellissard et al. [1989] Bellissard, J, B. Iochum, E. Scoppola, and D. Testard (1989), “Spectral properties of one dimensional quasi-crystals,” Commun. Math. Phys. 125 (3), 527–543.
- Benza et al. [1990] Benza, V G, M. Kolá, and M. K. Ali (1990), “Phase transition in the generalized Fibonacci quantum Ising models,” Phys. Rev. B 41, 9578–9580.
- Bera et al. [2017] Bera, Soumya, Thomas Martynec, Henning Schomerus, Fabian Heidrich-Meisner, and Jens H. Bardarson (2017), “One-particle density matrix characterization of many-body localization,” Ann. Phys. (Berl.) 529 (7), 1600356.
- Bercioux and Iñiguez [2018] Bercioux, Dario, and Ainhoa Iñiguez (2018), “Quantum fractals,” Nat. Phys. 15 (2), 111.
- Bercioux and Lucignano [2015] Bercioux, Dario, and Procolo Lucignano (2015), “Quantum transport in Rashba spin–orbit materials: a review,” Rep. Prog. Phys. 78 (10), 106001.
- Bhardwaj et al. [2021] Bhardwaj, Anshuman, Dzmitry Vaido, and Daniel E. Sheehy (2021), “Inflationary dynamics and particle production in a toroidal Bose-Einstein condensate,” Phys. Rev. A 103, 023322.
- Biddle et al. [2011] Biddle, J, D. J. Priour, B. Wang, and S. Das Sarma (2011), “Localization in one-dimensional lattices with non-nearest-neighbor hopping: Generalized Anderson and Aubry-André models,” Phys. Rev. B 83, 075105.
- Biesenthal et al. [2022] Biesenthal, Tobias, Lukas J. Maczewsky, Zhaoju Yang, Mark Kremer, Mordechai Segev, Alexander Szameit, and Matthias Heinrich (2022), “Fractal photonic topological insulators,” Science 376 (6597), 1114–1119.
- Billy et al. [2008] Billy, Juliette, Vincent Josse, Zhanchun Zuo, Alain Bernard, Ben Hambrecht, Pierre Lugan, David Clément, Laurent Sanchez-Palencia, Philippe Bouyer, and Alain Aspect (2008), “Direct observation of Anderson localization of matter waves in a controlled disorder,” Nature 453 (7197), 891–894.
- Boada et al. [2011] Boada, O, A Celi, J I Latorre, and M Lewenstein (2011), “Dirac equation for cold atoms in artificial curved spacetimes,” New J. Phys. 13 (3), 035002.
- Boada et al. [2012] Boada, O, A. Celi, J. I. Latorre, and M. Lewenstein (2012), “Quantum simulation of an extra dimension,” Phys. Rev. Lett. 108, 133001.
- Boada et al. [2015] Boada, Octavi, Alessio Celi, Javier Rodríguez-Laguna, José I Latorre, and Maciej Lewenstein (2015), “Quantum simulation of non-trivial topology,” New J. Phys. 17 (4), 045007.
- Boers et al. [2007] Boers, Dave J, Benjamin Goedeke, Dennis Hinrichs, and Martin Holthaus (2007), “Mobility edges in bichromatic optical lattices,” Phys. Rev. A 75, 063404.
- Boettcher et al. [2020] Boettcher, Igor, Przemyslaw Bienias, Ron Belyansky, Alicia J. Kollár, and Alexey V. Gorshkov (2020), “Quantum simulation of hyperbolic space with circuit quantum electrodynamics: From graphs to geometry,” Phys. Rev. A 102 (3), 032208.
- Boettcher et al. [2022] Boettcher, Igor, Alexey V. Gorshkov, Alicia J. Kollár, Joseph Maciejko, Steven Rayan, and Ronny Thomale (2022), “Crystallography of hyperbolic lattices,” Phys. Rev. B 105, 125118.
- Bouhiron et al. [2024] Bouhiron, Jean-Baptiste, Aurélien Fabre, Qi Liu, Quentin Redon, Nehal Mittal, Tanish Satoor, Raphael Lopes, and Sylvain Nascimbene (2024), “Realization of an atomic quantum Hall system in four dimensions,” Science 384 (6692), 223–227.
- Bukov et al. [2015] Bukov, Marin, Luca D’Alessio, and Anatoli Polkovnikov (2015), “Universal high-frequency behavior of periodically driven systems: from dynamical stabilization to Floquet engineering,” Adv. Phys. 64 (2), 139–226.
- Cai et al. [2013] Cai, Xiaoming, Li-Jun Lang, Shu Chen, and Yupeng Wang (2013), “Topological superconductor to anderson localization transition in one-dimensional incommensurate lattices,” Phys. Rev. Lett. 110, 176403.
- Canyellas et al. [2024] Canyellas, R, Chen Liu, R. Arouca, L. Eek, Guanyong Wang, Yin Yin, Dandan Guan, Yaoyi Li, Shiyong Wang, Hao Zheng, Canhua Liu, Jinfeng Jia, and C. Morais Smith (2024), “Topological edge and corner states in bismuth fractal nanostructures,” Nat. Phys. 10.1038/s41567-024-02551-8.
- Cao et al. [2018] Cao, Yuan, Valla Fatemi, Shiang Fang, Kenji Watanabe, Takashi Taniguchi, Efthimios Kaxiras, and Pablo Jarillo-Herrero (2018), “Unconventional superconductivity in magic-angle graphene superlattices,” Nature 556 (7699), 43–50.
- Caracanhas et al. [2022] Caracanhas, Mônica A, Pietro Massignan, and Alexander L. Fetter (2022), “Superfluid vortex dynamics on an ellipsoid and other surfaces of revolution,” Phys. Rev. A 105, 023307.
- Cardano et al. [2017] Cardano, Filippo, Alessio D’Errico, Alexandre Dauphin, Maria Maffei, Bruno Piccirillo, Corrado de Lisio, Giulio De Filippis, Vittorio Cataudella, Enrico Santamato, Lorenzo Marrucci, Maciej Lewenstein, and Pietro Massignan (2017), “Detection of Zak phases and topological invariants in a chiral quantum walk of twisted photons,” Nat. Commun. 8 (1), 15516.
- Cardano et al. [2016] Cardano, Filippo, Maria Maffei, Francesco Massa, Bruno Piccirillo, Corrado de Lisio, Giulio De Filippis, Vittorio Cataudella, Enrico Santamato, and Lorenzo Marrucci (2016), “Statistical moments of quantum-walk dynamics reveal topological quantum transitions,” Nat. Commun. 7 (1), 10.1038/ncomms11439.
- Carollo et al. [2022] Carollo, R A, D. C. Aveline, B. Rhyno, S. Vishveshwara, C. Lannert, J. D. Murphree, E. R. Elliott, J. R. Williams, R. J. Thompson, and N. Lundblad (2022), “Observation of ultracold atomic bubbles in orbital microgravity,” Nature 606 (7913), 281–286.
- Carusotto and Ciuti [2013] Carusotto, Iacopo, and Cristiano Ciuti (2013), “Quantum fluids of light,” Rev. Mod. Phys. 85, 299–366.
- Carusotto et al. [2008] Carusotto, Iacopo, Serena Fagnocchi, Alessio Recati, Roberto Balbinot, and Alessandro Fabbri (2008), “Numerical observation of Hawking radiation from acoustic black holes in atomic Bose–Einstein condensates,” New J. Phys. 10 (10), 103001.
- Chalopin et al. [2020] Chalopin, Thomas, Tanish Satoor, Alexandre Evrard, Vasiliy Makhalov, Jean Dalibard, Raphael Lopes, and Sylvain Nascimbene (2020), “Probing chiral edge dynamics and bulk topology of a synthetic Hall system,” Nat. Phys. 16 (10), 1017–1021, 2001.01664 .
- Chen et al. [2023] Chen, Anffany, Hauke Brand, Tobias Helbig, Tobias Hofmann, Stefan Imhof, Alexander Fritzsche, Tobias Kießling, Alexander Stegmaier, Lavi K. Upreti, Titus Neupert, Tomáš Bzdušek, Martin Greiter, Ronny Thomale, and Igor Boettcher (2023), “Hyperbolic matter in electrical circuits with tunable complex phases,” Nat. Commun. 14 (1), 622.
- Chen et al. [2024] Chen, Anffany, Joseph Maciejko, and Igor Boettcher (2024), “Anderson localization transition in disordered hyperbolic lattices,” Phys. Rev. Lett. 133, 066101.
- Chen et al. [2020] Chen, Rui, Chui-Zhen Chen, Jin-Hua Gao, Bin Zhou, and Dong-Hui Xu (2020), “Higher-order topological insulators in quasicrystals,” Phys. Rev. Lett. 124, 036803.
- Cheng et al. [2019] Cheng, Peng, Philipp W. Klein, Kirill Plekhanov, Klaus Sengstock, Monika Aidelsburger, Christof Weitenberg, and Karyn Le Hur (2019), “Topological proximity effects in a Haldane graphene bilayer system,” Phys. Rev. B 100, 081107.
- Chiaracane et al. [2021] Chiaracane, Cecilia, Francesca Pietracaprina, Archak Purkayastha, and John Goold (2021), “Quantum dynamics in the interacting Fibonacci chain,” Phys. Rev. B 103, 184205.
- Chin et al. [2010] Chin, Cheng, Rudolf Grimm, Paul Julienne, and Eite Tiesinga (2010), “Feshbach resonances in ultracold gases,” Rev. Mod. Phys. 82, 1225–1286.
- Chiu et al. [2016] Chiu, Ching-Kai, Jeffrey C. Y. Teo, Andreas P. Schnyder, and Shinsei Ryu (2016), “Classification of topological quantum matter with symmetries,” Rev. Mod. Phys. 88, 035005.
- Choi et al. [2016] Choi, Jae-yoon, Sebastian Hild, Johannes Zeiher, Peter Schauß, Antonio Rubio-Abadal, Tarik Yefsah, Vedika Khemani, David A. Huse, Immanuel Bloch, and Christian Gross (2016), “Exploring the many-body localization transition in two dimensions,” Science 352 (6293), 1547–1552.
- Clark et al. [2020] Clark, Logan W, Nathan Schine, Claire Baum, Ningyuan Jia, and Jonathan Simon (2020), “Observation of Laughlin states made of light,” Nature 582 (7810), 41–45.
- Collins et al. [2017] Collins, Laura C, Thomas G. Witte, Rochelle Silverman, David B. Green, and Kenjiro K. Gomes (2017), “Imaging quasiperiodic electronic states in a synthetic Penrose tiling,” Nat. Commun. 8 (1), 15961.
- Corcovilos and Mittal [2019] Corcovilos, Theodore A, and Jahnavee Mittal (2019), “Two-dimensional optical quasicrystal potentials for ultracold atom experiments,” Appl. Opt. 58 (9), 2256–2263.
- Coxeter [1973] Coxeter, H S M (1973), Regular Polytopes, Dover Books on Mathematics (Dover Publications, Mineola, NY).
- Darázs et al. [2014] Darázs, Zoltán, Anastasiia Anishchenko, Tamás Kiss, Alexander Blumen, and Oliver Mülken (2014), “Transport properties of continuous-time quantum walks on sierpinski fractals,” Phys. Rev. E 90, 032113.
- Dareau et al. [2017] Dareau, A, E. Levy, M. Bosch Aguilera, R. Bouganne, E. Akkermans, F. Gerbier, and J. Beugnon (2017), “Revealing the topology of quasicrystals with a diffraction experiment,” Phys. Rev. Lett. 119, 215304.
- Das Sarma et al. [1988] Das Sarma, S, Song He, and X. C. Xie (1988), “Mobility edge in a model one-dimensional potential,” Phys. Rev. Lett. 61, 2144–2147.
- Davies [1975] Davies, P C W (1975), “Scalar production in Schwarzschild and Rindler metrics,” J. Phys. A: Math. Gen. 8 (4), 609.
- De Beule et al. [2021] De Beule, Christophe, Solofo Groenendijk, Tobias Meng, and Thomas L. Schmidt (2021), “Artificial event horizons in Weyl semimetal heterostructures and their non-equilibrium signatures,” SciPost Phys. 11, 095.
- DeGottardi et al. [2013a] DeGottardi, Wade, Diptiman Sen, and Smitha Vishveshwara (2013a), “Majorana Fermions in Superconducting 1D Systems Having Periodic, Quasiperiodic, and Disordered Potentials,” Phys. Rev. Lett. 110, 146404.
- DeGottardi et al. [2013b] DeGottardi, Wade, Manisha Thakurathi, Smitha Vishveshwara, and Diptiman Sen (2013b), “Majorana fermions in superconducting wires: Effects of long-range hopping, broken time-reversal symmetry, and potential landscapes,” Phys. Rev. B 88, 165111.
- Deng et al. [2019] Deng, X, S. Ray, S. Sinha, G. V. Shlyapnikov, and L. Santos (2019), “One-dimensional quasicrystals with power-law hopping,” Phys. Rev. Lett. 123, 025301.
- Di Carli et al. [2020] Di Carli, Andrea, Grant Henderson, Stuart Flannigan, Craig D. Colquhoun, Matthew Mitchell, Gian-Luca Oppo, Andrew J. Daley, Stefan Kuhr, and Elmar Haller (2020), “Collisionally Inhomogeneous Bose-Einstein Condensates with a Linear Interaction Gradient,” Phys. Rev. Lett. 125, 183602.
- Doggen and Mirlin [2019] Doggen, Elmer V H, and Alexander D. Mirlin (2019), “Many-body delocalization dynamics in long Aubry-André quasiperiodic chains,” Phys. Rev. B 100, 104203.
- Domany et al. [1983] Domany, Eytan, Shlomo Alexander, David Bensimon, and Leo P. Kadanoff (1983), “Solutions to the schrödinger equation on some fractal lattices,” Phys. Rev. B 28, 3110–3123.
- Dutt et al. [2020] Dutt, Avik, Qian Lin, Luqi Yuan, Momchil Minkov, Meng Xiao, and Shanhui Fan (2020), “A single photonic cavity with two independent physical synthetic dimensions,” Science 367 (6473), 59–64.
- Dutt et al. [2019] Dutt, Avik, Momchil Minkov, Qian Lin, Luqi Yuan, David A. B. Miller, and Shanhui Fan (2019), “Experimental band structure spectroscopy along a synthetic dimension,” Nat. Commun. 10 (1), 3122.
- Ebadi et al. [2022] Ebadi, S, A. Keesling, M. Cain, T. T. Wang, H. Levine, D. Bluvstein, G. Semeghini, A. Omran, J.-G. Liu, R. Samajdar, X.-Z. Luo, B. Nash, X. Gao, B. Barak, E. Farhi, S. Sachdev, N. Gemelke, L. Zhou, S. Choi, H. Pichler, S.-T. Wang, M. Greiner, V. Vuletić, and M. D. Lukin (2022), “Quantum optimization of maximum independent set using Rydberg atom arrays,” Science 376 (6598), 1209–1215.
- Eckardt [2017] Eckardt, André (2017), “Colloquium: Atomic quantum gases in periodically driven optical lattices,” Rev. Mod. Phys. 89, 011004.
- Eckel et al. [2018] Eckel, S, A. Kumar, T. Jacobson, I. B. Spielman, and G. K. Campbell (2018), “A Rapidly Expanding Bose-Einstein Condensate: An Expanding Universe in the Lab,” Phys. Rev. X 8, 021021.
- Ehrhardt et al. [2024] Ehrhardt, Max, Christoph Dittel, Matthias Heinrich, and Alexander Szameit (2024), “Topological Hong-Ou-Mandel interference,” Science 384 (6702), 1340–1344.
- Ehrhardt et al. [2021] Ehrhardt, Max, Robert Keil, Lukas J. Maczewsky, Christoph Dittel, Matthias Heinrich, and Alexander Szameit (2021), “Exploring complex graphs using three-dimensional quantum walks of correlated photons,” Sci. Adv. 7 (9), eabc5266.
- Eiguren et al. [2002] Eiguren, A, B. Hellsing, F. Reinert, G. Nicolay, E. V. Chulkov, V. M. Silkin, S. Hüfner, and P. M. Echenique (2002), “Role of bulk and surface phonons in the decay of metal surface states,” Phys. Rev. Lett. 88, 066805.
- Else et al. [2021] Else, Dominic V, Sheng-Jie Huang, Abhinav Prem, and Andrey Gromov (2021), “Quantum many-body topology of quasicrystals,” Phys. Rev. X 11, 041051.
- Endres et al. [2016] Endres, Manuel, Hannes Bernien, Alexander Keesling, Harry Levine, Eric R. Anschuetz, Alexandre Krajenbrink, Crystal Senko, Vladan Vuletic, Markus Greiner, and Mikhail D. Lukin (2016), “Atom-by-atom assembly of defect-free one-dimensional cold atom arrays,” Science 354 (6315), 1024–1027.
- Evers and Mirlin [2008] Evers, Ferdinand, and Alexander D. Mirlin (2008), “Anderson transitions,” Rev. Mod. Phys. 80, 1355–1417.
- Feynman [1982] Feynman, Richard P (1982), “Simulating physics with computers,” Int. J. Theor. Phys. 21 (6), 467–488.
- Grossi e Fonseca et al. [2023] Grossi e Fonseca, André, Thomas Christensen, John D. Joannopoulos, and Marin Soljačić (2023), “Quasicrystalline weyl points and dense fermi-bragg arcs,” Phys. Rev. B 108, L121109.
- Fonseca et al. [2024] Fonseca, André Grossi, Sachin Vaidya, Thomas Christensen, Mikael C. Rechtsman, Taylor L. Hughes, and Marin Soljačić (2024), “Weyl points on nonorientable manifolds,” Phys. Rev. Lett. 132, 266601.
- Fournée et al. [2014] Fournée, Vincent, Émilie Gaudry, Julian Ledieu, Marie-Cécile de Weerd, Dongmei Wu, and Thomas Lograsso (2014), “Self-Organized Molecular Films with Long-Range Quasiperiodic Order,” ACS Nano 8 (4), 3646–3653.
- Fraxanet et al. [2022] Fraxanet, Joana, Utso Bhattacharya, Tobias Grass, Maciej Lewenstein, and Alexandre Dauphin (2022), “Localization and multifractal properties of the long-range Kitaev chain in the presence of an Aubry-André-Harper modulation,” Phys. Rev. B 106, 024204.
- Fraxanet et al. [2021] Fraxanet, Joana, Utso Bhattacharya, Tobias Grass, Debraj Rakshit, Maciej Lewenstein, and Alexandre Dauphin (2021), “Topological properties of the long-range Kitaev chain with Aubry-André-Harper modulation,” Phys. Rev. Res. 3, 013148.
- Freedman et al. [2006] Freedman, Barak, Guy Bartal, Mordechai Segev, Ron Lifshitz, Demetrios N. Christodoulides, and Jason W. Fleischer (2006), “Wave and defect dynamics in nonlinear photonic quasicrystals,” Nature 440 (7088), 1166–1169.
- Freeney et al. [2020] Freeney, S E, J. J. van den Broeke, A. J. J. Harsveld van der Veen, I. Swart, and C. Morais Smith (2020), “Edge-Dependent Topology in Kekulé Lattices,” Phys. Rev. Lett. 124 (23), 236404.
- Freeney et al. [2022] Freeney, Saoirsé E, Marlou R. Slot, Thomas S. Gardenier, Ingmar Swart, and Daniel Vanmaekelbergh (2022), “Electronic quantum materials simulated with artificial model lattices,” ACS Nanosci. Au 2 (3), 198–224.
- Fremling et al. [2020] Fremling, Mikael, Michal van Hooft, Cristiane Morais Smith, and Lars Fritz (2020), “Existence of robust edge currents in Sierpiński fractals,” Phys. Rev. Res. 2, 013044.
- Fukui et al. [2005] Fukui, Takahiro, Yasuhiro Hatsugai, and Hiroshi Suzuki (2005), “Chern Numbers in Discretized Brillouin Zone: Efficient Method of Computing (Spin) Hall Conductances,” J. Phys. Soc. Jpn. 74 (6), 1674–1677.
- Fulling [1973] Fulling, Stephen A (1973), “Nonuniqueness of Canonical Field Quantization in Riemannian Space-Time,” Phys. Rev. D 7, 2850–2862.
- Ganeshan et al. [2015] Ganeshan, Sriram, J. H. Pixley, and S. Das Sarma (2015), “Nearest neighbor tight binding models with an exact mobility edge in one dimension,” Phys. Rev. Lett. 114, 146601.
- Garay et al. [2000] Garay, L J, J. R. Anglin, J. I. Cirac, and P. Zoller (2000), “Sonic Analog of Gravitational Black Holes in Bose-Einstein Condensates,” Phys. Rev. Lett. 85, 4643–4647.
- Garraway and Perrin [2016] Garraway, Barry M, and Hélène Perrin (2016), “Recent developments in trapping and manipulation of atoms with adiabatic potentials,” J. Phys. B 49 (17), 172001.
- Gautier et al. [2021] Gautier, Ronan, Hepeng Yao, and Laurent Sanchez-Palencia (2021), “Strongly interacting bosons in a two-dimensional quasicrystal lattice,” Phys. Rev. Lett. 126, 110401.
- Goblot et al. [2020] Goblot, V, A. Štrkalj, N. Pernet, J. L. Lado, C. Dorow, A. Lemaître, L. Le Gratiet, A. Harouri, I. Sagnes, S. Ravets, A. Amo, J. Bloch, and O. Zilberberg (2020), “Emergence of criticality through a cascade of delocalization transitions in quasiperiodic chains,” Nat. Phys. 16 (8), 832–836.
- Gomes et al. [2012] Gomes, Kenjiro K, Warren Mar, Wonhee Ko, Francisco Guinea, and Hari C kran (2012), “Designer Dirac fermions and topological phases in molecular graphene,” Nature 483 (7389), 306 – 310.
- Gottlob and Schneider [2023] Gottlob, E, and U. Schneider (2023), “Hubbard models for quasicrystalline potentials,” Phys. Rev. B 107, 144202.
- Graß et al. [2015] Graß, Tobias, Christine Muschik, Alessio Celi, Ravindra W. Chhajlany, and Maciej Lewenstein (2015), “Synthetic magnetic fluxes and topological order in one-dimensional spin systems,” Phys. Rev. A 91, 063612.
- Haake [2001] Haake, F (2001), Quantum Signatures of Chaos, Physics and astronomy online library (Springer).
- Halsey et al. [1986] Halsey, Thomas C, Mogens H. Jensen, Leo P. Kadanoff, Itamar Procaccia, and Boris I. Shraiman (1986), “Fractal measures and their singularities: The characterization of strange sets,” Phys. Rev. A 33, 1141–1151.
- Hao et al. [2024] Hao, Chen-Yue, Zhen Zhan, Pierre A. Pantaleón, Jia-Qi He, Ya-Xin Zhao, Kenji Watanabe, Takashi Taniguchi, Francisco Guinea, and Lin He (2024), “Robust flat bands in twisted trilayer graphene moiré quasicrystals,” Nat. Commun. 15 (1), 8437.
- Harper [1955] Harper, P G (1955), “Single band motion of conduction electrons in a uniform magnetic field,” Proc. Phys. Soc. 68 (10), 874.
- Hars et al. [2018] Hars, SS, H.R. Sharma, J.A. Smerdon, S. Coates, K. Nozawa, A.P. Tsai, and R. McGrath (2018), “Growth of a bismuth thin film on the five-fold surface of the icosahedral Ag-In-Yb quasicrystal,” Surf. Sci. 678, 222–227.
- Hatsugai [1993] Hatsugai, Yasuhiro (1993), “Chern number and edge states in the integer quantum Hall effect,” Phys. Rev. Lett. 71, 3697–3700.
- Hawking [1974] Hawking, S W (1974), “Black hole explosions?” Nature 248 (5443), 30.
- Hendrickson et al. [2008] Hendrickson, J, B. C. Richards, J. Sweet, G. Khitrova, A. N. Poddubny, E. L. Ivchenko, M. Wegener, and H. M. Gibbs (2008), “Excitonic polaritons in Fibonacci quasicrystals,” Opt. Express 16 (20), 15382–15387.
- Hida [1999] Hida, Kazuo (1999), “Density Matrix Renormalization Group Study of the S= 1/2 Antiferromagnetic Heisenberg Chains with Quasiperiodic Exchange Modulation,” J. Phys. Soc. Jpn. 68 (10), 3177–3180.
- Hida [2004] Hida, Kazuo (2004), “New universality class in spin-one-half fibonacci heisenberg chains,” Phys. Rev. Lett. 93, 037205.
- Hiramoto and Kohmoto [1989] Hiramoto, Hisashi, and Mahito Kohmoto (1989), “Scaling analysis of quasiperiodic systems: Generalized Harper model,” Phys. Rev. B 40, 8225–8234.
- Hofstadter [1976] Hofstadter, Douglas R (1976), “Energy levels and wave functions of bloch electrons in rational and irrational magnetic fields,” Phys. Rev. B 14, 2239–2249.
- Hu et al. [2019] Hu, Jiazhong, Lei Feng, Zhendong Zhang, and Cheng Chin (2019), “Quantum simulation of Unruh radiation,” Nat. Phys. 15 (8), 785–789.
- Huang et al. [2024] Huang, Lei, Lu He, Weixuan Zhang, Huizhen Zhang, Dongning Liu, Xue Feng, Fang Liu, Kaiyu Cui, Yidong Huang, Wei Zhang, and Xiangdong Zhang (2024), “Hyperbolic photonic topological insulators,” Nat. Commun. 15 (1), 1–9.
- Huang et al. [2020] Huang, Xueqin, Weiyin Deng, Feng Li, Jiuyang Lu, and Zhengyou Liu (2020), “Ideal Type-II Weyl Phase and Topological Transition in Phononic Crystals,” Phys. Rev. Lett. 124, 206802.
- Iglói et al. [1997] Iglói, Ferenc, Loïc Turban, Dragi Karevski, and Ferenc Szalma (1997), “Exact renormalization-group study of aperiodic Ising quantum chains and directed walks,” Phys. Rev. B 56, 11031–11050.
- Iglói et al. [2007] Iglói, F, R. Juhász, and Z. Zimborás (2007), “Entanglement entropy of aperiodic quantum spin chains,” EPL 79 (3), 37001.
- Iliasov et al. [2020] Iliasov, Askar A, Mikhail I. Katsnelson, and Shengjun Yuan (2020), “Hall conductivity of a Sierpiński carpet,” Phys. Rev. B 101, 045413.
- Isenhower et al. [2010] Isenhower, L, E. Urban, X. L. Zhang, A. T. Gill, T. Henage, T. A. Johnson, T. G. Walker, and M. Saffman (2010), “Demonstration of a neutral atom controlled-not quantum gate,” Phys. Rev. Lett. 104, 010503.
- Iyer et al. [2013] Iyer, Shankar, Vadim Oganesyan, Gil Refael, and David A. Huse (2013), “Many-body localization in a quasiperiodic system,” Phys. Rev. B 87, 134202.
- Jacqmin et al. [2014] Jacqmin, T, I. Carusotto, I. Sagnes, M. Abbarchi, D. D. Solnyshkov, G. Malpuech, E. Galopin, A. Lemaître, J. Bloch, and A. Amo (2014), “Direct observation of dirac cones and a flatband in a honeycomb lattice for polaritons,” Phys. Rev. Lett. 112, 116402.
- Jagannathan [2021] Jagannathan, Anuradha (2021), “The Fibonacci quasicrystal: Case study of hidden dimensions and multifractality,” Rev. Mod. Phys. 93, 045001.
- Jamadi et al. [2020] Jamadi, Omar, Elena Rozas, Grazia Salerno, Marijana Milićević, Tomoki Ozawa, Isabelle Sagnes, Aristide Lemaître, Luc Le Gratiet, Abdelmounaim Harouri, Iacopo Carusotto, Jacqueline Bloch, and Alberto Amo (2020), “Direct observation of photonic Landau levels and helical edge states in strained honeycomb lattices,” Light Sci. Appl. 9 (1), 144.
- Johnstone et al. [2019] Johnstone, Dean, Patrik Öhberg, and Callum W. Duncan (2019), “Mean-field phases of an ultracold gas in a quasicrystalline potential,” Phys. Rev. A 100, 053609.
- Kempkes et al. [2019a] Kempkes, S N, M. R. Slot, J. J. van den Broeke, P. Capiod, W. A. Benalcazar, D. Vanmaekelbergh, D. Bercioux, I. Swart, and C. Morais Smith (2019a), “Robust zero-energy modes in an electronic higher-order topological insulator,” Nat. Mater. 18 (12), 1292–1297.
- Kempkes et al. [2019b] Kempkes, S N, M. R. Slot, S. E. Freeney, S. J. M. Zevenhuizen, D. Vanmaekelbergh, I. Swart, and C. Morais Smith (2019b), “Design and characterization of electrons in a fractal geometry,” Nat. Phys. 15 (2), 127–131.
- Kennes et al. [2021] Kennes, Dante M, Martin Claassen, Lede Xian, Antoine Georges, Andrew J. Millis, James Hone, Cory R. Dean, D. N. Basov, Abhay N. Pasupathy, and Angel Rubio (2021), “Moiré heterostructures as a condensed-matter quantum simulator,” Nature Physics 17 (2), 155–163.
- Khajetoorians et al. [2019] Khajetoorians, Alexander A, Daniel Wegner, Alexander F. Otte, and Ingmar Swart (2019), “Creating designer quantum states of matter atom-by-atom,” Nat. Rev. Phys. 1 (12), 703–715.
- Khemani et al. [2017] Khemani, Vedika, D. N. Sheng, and David A. Huse (2017), “Two universality classes for the many-body localization transition,” Phys. Rev. Lett. 119, 075702.
- Khveshchenko [2015] Khveshchenko, D V (2015), “Analogue holographic correspondence in optical metamaterials,” Europhysics Letters 109 (6), 61001.
- Kitaev [2001] Kitaev, A Yu (2001), “Unpaired Majorana fermions in quantum wires,” Phys.-Uspekhi 44 (10S), 131.
- Klembt et al. [2018] Klembt, S, T H Harder, O A Egorov, K Winkler, R Ge, M A Bandres, M Emmerling, L Worschech, T C H Liew, M Segev, C Schneider, and S Höfling (2018), “Exciton-polariton topological insulator,” Nature 562 (7728), 552–556.
- Kohlert et al. [2019] Kohlert, Thomas, Sebastian Scherg, Xiao Li, Henrik P. Lüschen, Sankar Das Sarma, Immanuel Bloch, and Monika Aidelsburger (2019), “Observation of Many-Body Localization in a One-Dimensional System with a Single-Particle Mobility Edge,” Phys. Rev. Lett. 122, 170403.
- Kohmoto et al. [1983] Kohmoto, Mahito, Leo P. Kadanoff, and Chao Tang (1983), “Localization problem in one dimension: Mapping and escape,” Phys. Rev. Lett. 50, 1870–1872.
- Kollár et al. [2019] Kollár, Alicia J, Mattias Fitzpatrick, and Andrew A. Houck (2019), “Hyperbolic lattices in circuit quantum electrodynamics,” Nature 571 (7763), 45–50.
- Kolobov et al. [2021] Kolobov, Victor I, Katrine Golubkov, Juan Ramón Muñoz de Nova, and Jeff Steinhauer (2021), “Observation of stationary spontaneous Hawking radiation and the time evolution of an analogue black hole,” Nat. Phys. 17 (3), 362–367.
- Kosch et al. [2022] Kosch, M N, L. Asteria, H. P. Zahn, K. Sengstock, and C. Weitenberg (2022), “Multifrequency optical lattice for dynamic lattice-geometry control,” Phys. Rev. Res. 4, 043083.
- Kosior et al. [2018] Kosior, Arkadiusz, Maciej Lewenstein, and Alessio Celi (2018), “Unruh effect for interacting particles with ultracold atoms,” SciPost Phys. 5, 061.
- Kosior and Sacha [2017] Kosior, Arkadiusz, and Krzysztof Sacha (2017), “Localization in random fractal lattices,” Phys. Rev. B 95, 104206.
- Kraus et al. [2012] Kraus, Yaacov E, Yoav Lahini, Zohar Ringel, Mor Verbin, and Oded Zilberberg (2012), “Topological states and adiabatic pumping in quasicrystals,” Phys. Rev. Lett. 109, 106402.
- Kraus et al. [2013] Kraus, Yaacov E, Zohar Ringel, and Oded Zilberberg (2013), “Four-Dimensional Quantum Hall Effect in a Two-Dimensional Quasicrystal,” Phys. Rev. Lett. 111, 226401.
- Kraus and Zilberberg [2012] Kraus, Yaacov E, and Oded Zilberberg (2012), “Topological Equivalence between the Fibonacci Quasicrystal and the Harper Model,” Phys. Rev. Lett. 109, 116404.
- Küster et al. [2022] Küster, Felix, Sascha Brinker, Richard Hess, Daniel Loss, Stuart S. P. Parkin, Jelena Klinovaja, Samir Lounis, and Paolo Sessi (2022), “Non-Majorana modes in diluted spin chains proximitized to a superconductor,” Proc. Natl. Acad. Sci. U.S.A. 119 (42), e2210589119.
- Lahini et al. [2009] Lahini, Y, R. Pugatch, F. Pozzi, M. Sorel, R. Morandotti, N. Davidson, and Y. Silberberg (2009), “Observation of a localization transition in quasiperiodic photonic lattices,” Phys. Rev. Lett. 103, 013901.
- Lee et al. [2018] Lee, Ching Hua, Stefan Imhof, Christian Berger, Florian Bayer, Johannes Brehm, Laurens W. Molenkamp, Tobias Kiessling, and Ronny Thomale (2018), “Topolectrical Circuits,” Commun. Phys. 1 (1), 39.
- Lee et al. [2022] Lee, Ki Young, Kwang Wook Yoo, Sangmo Cheon, Won-Jae Joo, Jae Woong Yoon, and Seok Ho Song (2022), “Synthetic Topological Nodal Phase in Bilayer Resonant Gratings,” Phys. Rev. Lett. 128, 053002.
- Lee et al. [2017] Lee, Mac, Thomas R. Look, S. P. Lim, and D. N. Sheng (2017), “Many-body localization in spin chain systems with quasiperiodic fields,” Phys. Rev. B 96, 075146.
- Lenggenhager et al. [2022] Lenggenhager, Patrick M, Alexander Stegmaier, Lavi K. Upreti, Tobias Hofmann, Tobias Helbig, Achim Vollhardt, Martin Greiter, Ching Hua Lee, Stefan Imhof, Hauke Brand, Tobias Kießling, Igor Boettcher, Titus Neupert, Ronny Thomale, and Tomáš Bzdušek (2022), “Simulating hyperbolic space on a circuit board,” Nat. Commun. 13 (1), 4373.
- Lev et al. [2017] Lev, Yevgeny Bar, Dante M. Kennes, Christian Klöckner, David R. Reichman, and Christoph Karrasch (2017), “Transport in quasiperiodic interacting systems: From superdiffusion to subdiffusion,” EPL 119 (3), 37003.
- Levine and Steinhardt [1984] Levine, Dov, and Paul Joseph Steinhardt (1984), “Quasicrystals: A new class of ordered structures,” Phys. Rev. Lett. 53, 2477–2480.
- Li et al. [2022a] Li, Chuan-Hsun, Yangqian Yan, Shih-Wen Feng, Sayan Choudhury, David B. Blasing, Qi Zhou, and Yong P. Chen (2022a), “Bose-Einstein Condensate on a Synthetic Topological Hall Cylinder,” PRX Quantum 3, 010316.
- Li et al. [2023a] Li, Hao, Yong-Yi Wang, Yun-Hao Shi, Kaixuan Huang, Xiaohui Song, Gui-Han Liang, Zheng-Yang Mei, Bozhen Zhou, He Zhang, Jia-Chi Zhang, Shu Chen, S. P. Zhao, Ye Tian, Zhan-Ying Yang, Zhongcheng Xiang, Kai Xu, Dongning Zheng, and Heng Fan (2023a), “Observation of critical phase transition in a generalized aubry-andré-harper model with superconducting circuits,” npj Quantum Information 9 (1), 40.
- Li et al. [2023b] Li, Meng, Chu Li, Linyu Yan, Qiang Li, Qihuang Gong, and Yan Li (2023b), “Fractal photonic anomalous Floquet topological insulators to generate multiple quantum chiral edge states,” Light Sci. Appl. 12 (1), 262.
- Li et al. [2017] Li, Xiao, Xiaopeng Li, and S. Das Sarma (2017), “Mobility edges in one-dimensional bichromatic incommensurate potentials,” Phys. Rev. B 96, 085119.
- Li et al. [2022b] Li, Xiaoxia, Qili Li, Tongzhou Ji, Ruige Yan, Wenlin Fan, Bingfeng Miao, Liang Sun, Gong Chen, Weiyi Zhang, and Haifeng Ding (2022b), “Lieb Lattices Formed by Real Atoms on Ag(111) and Their Lattice Constant-Dependent Electronic Properties,” Chin. Phys. Lett. 39 (5), 057301.
- Li et al. [2022c] Li, Xikun, Mani Chandra Jha, and Anne E. B. Nielsen (2022c), “Laughlin topology on fractal lattices without area law entanglement,” Phys. Rev. B 105, 085152.
- Li et al. [2024] Li, Yanxing, Fan Zhang, Viet-Anh Ha, Yu-Chuan Lin, Chengye Dong, Qiang Gao, Zhida Liu, Xiaohui Liu, Sae Hee Ryu, Hyunsue Kim, Chris Jozwiak, Aaron Bostwick, Kenji Watanabe, Takashi Taniguchi, Bishoy Kousa, Xiaoqin Li, Eli Rotenberg, Eslam Khalaf, Joshua A. Robinson, Feliciano Giustino, and Chih-Kang Shih (2024), “Tuning commensurability in twisted van der Waals bilayers,” Nature 625 (7995), 494–499.
- Lifshitz [2003] Lifshitz, Ron (2003), “Quasicrystals: A matter of definition,” Found. Phys. 33 (12), 1703–1711.
- Liu et al. [2021] Liu, Chen, Yinong Zhou, Guanyong Wang, Yin Yin, Can Li, Haili Huang, Dandan Guan, Yaoyi Li, Shiyong Wang, Hao Zheng, Canhua Liu, Yong Han, James W. Evans, Feng Liu, and Jinfeng Jia (2021), “Sierpiński Structure and Electronic Topology in Bi Thin Films on InSb(111)B Surfaces,” Phys. Rev. Lett. 126 (17), 176102.
- Lohse et al. [2016] Lohse, M, C. Schweizer, O. Zilberberg, M. Aidelsburger, and I. Bloch (2016), “A Thouless quantum pump with ultracold bosonic atoms in an optical superlattice,” Nat. Phys. 12 (4), 350–354.
- Lohse et al. [2018] Lohse, Michael, Christian Schweizer, Hannah M. Price, Oded Zilberberg, and Immanuel Bloch (2018), “Exploring 4D quantum Hall physics with a 2D topological charge pump,” Nature 553 (7686), 55–58.
- Luck [1993] Luck, J M (1993), “Critical behavior of the aperiodic quantum ising chain in a transverse magnetic field,” J. Stat. Phys. 72 (3), 417–458.
- Lukin et al. [2019] Lukin, Alexander, Matthew Rispoli, Robert Schittko, M. Eric Tai, Adam M. Kaufman, Soonwon Choi, Vedika Khemani, Julian Léonard, and Markus Greiner (2019), “Probing entanglement in a many-body localized system,” Science 364 (6437), 256–260.
- Lundblad et al. [2023] Lundblad, Nathan, David C Aveline, Antun Balaž, Elliot Bentine, Nicholas P Bigelow, Patrick Boegel, Maxim A Efremov, Naceur Gaaloul, Matthias Meister, Maxim Olshanii, Carlos A R Sá de Melo, Andrea Tononi, Smitha Vishveshwara, Angela C White, Alexander Wolf, and Barry M Garraway (2023), “Perspective on quantum bubbles in microgravity,” Quantum Science and Technology 8 (2), 024003.
- Lüschen et al. [2017] Lüschen, Henrik P, Pranjal Bordia, Sebastian Scherg, Fabien Alet, Ehud Altman, Ulrich Schneider, and Immanuel Bloch (2017), “Observation of slow dynamics near the many-body localization transition in one-dimensional quasiperiodic systems,” Phys. Rev. Lett. 119, 260401.
- Lüschen et al. [2018] Lüschen, Henrik P, Sebastian Scherg, Thomas Kohlert, Michael Schreiber, Pranjal Bordia, Xiao Li, S. Das Sarma, and Immanuel Bloch (2018), “Single-particle mobility edge in a one-dimensional quasiperiodic optical lattice,” Phys. Rev. Lett. 120, 160404.
- Lustig et al. [2022] Lustig, Eran, Lukas J. Maczewsky, Julius Beck, Tobias Biesenthal, Matthias Heinrich, Zhaoju Yang, Yonatan Plotnik, Alexander Szameit, and Mordechai Segev (2022), “Photonic topological insulator induced by a dislocation in three dimensions,” Nature 609 (7929), 931–935.
- Macé et al. [2019] Macé, Nicolas, Nicolas Laflorencie, and Fabien Alet (2019), “Many-body localization in a quasiperiodic Fibonacci chain,” SciPost Phys. 6, 050.
- Madsen et al. [2013] Madsen, Kevin A, Emil J. Bergholtz, and Piet W. Brouwer (2013), “Topological equivalence of crystal and quasicrystal band structures,” Phys. Rev. B 88, 125118.
- Maldacena [1999] Maldacena, Juan (1999), “The Large-N Limit of Superconformal Field Theories and Supergravity,” Int. J. Theor. Phys. 38 (4), 1113–1133.
- Mancini et al. [2015] Mancini, M, G. Pagano, G. Cappellini, L. Livi, M. Rider, J. Catani, C. Sias, P. Zoller, M. Inguscio, M. Dalmonte, and L. Fallani (2015), “Observation of chiral edge states with neutral fermions in synthetic Hall ribbons,” Science 349 (6255), 1510–1513.
- Mandelbrot [1982] Mandelbrot, Benoit B (1982), The fractal geometry of nature (W. H. Freeman and Co., San Francisco, Calif.).
- Manna et al. [2022] Manna, Sourav, Callum W. Duncan, Carrie A. Weidner, Jacob F. Sherson, and Anne E. B. Nielsen (2022), “Anyon braiding on a fractal lattice with a local Hamiltonian,” Phys. Rev. A 105, L021302.
- Manna et al. [2024] Manna, Sourav, Błażej Jaworowski, and Anne E B Nielsen (2024), “Many-body localization on finite generation fractal lattices,” J. Stat. Mech.: Theory Exp. 2024 (5), 053301.
- Manna et al. [2020] Manna, Sourav, Biplab Pal, Wei Wang, and Anne E. B. Nielsen (2020), “Anyons and fractional quantum Hall effect in fractal dimensions,” Phys. Rev. Res. 2, 023401.
- Matsuda et al. [2014] Matsuda, Fuyuki, Masaki Tezuka, and Norio Kawakami (2014), “Topological Properties of Ultracold Bosons in One-Dimensional Quasiperiodic Optical Lattice,” J. Phys. Soc. Jpn. 83 (8), 083707.
- Merlin et al. [1985] Merlin, R, K. Bajema, Roy Clarke, F. Y. Juang, and P. K. Bhattacharya (1985), “Quasiperiodic GaAs-AlAs Heterostructures,” Phys. Rev. Lett. 55, 1768–1770.
- Mochol-Grzelak et al. [2018] Mochol-Grzelak, M, A Dauphin, A Celi, and M Lewenstein (2018), “Efficient algorithm to compute the second Chern number in four dimensional systems,” Quantum Sci. Technol. 4 (1), 014009.
- Myerson-Jain et al. [2022] Myerson-Jain, Nayan E, Stephen Yan, David Weld, and Cenke Xu (2022), “Construction of Fractal Order and Phase Transition with Rydberg Atoms,” Phys. Rev. Lett. 128, 017601.
- Nadj-Perge et al. [2014] Nadj-Perge, Stevan, Ilya K. Drozdov, Jian Li, Hua Chen, Sangjun Jeon, Jungpil Seo, Allan H. MacDonald, B. Andrei Bernevig, and Ali Yazdani (2014), “Observation of Majorana fermions in ferromagnetic atomic chains on a superconductor,” Science 346 (6209), 602–607.
- Nakajima et al. [2016] Nakajima, Shuta, Takafumi Tomita, Shintaro Taie, Tomohiro Ichinose, Hideki Ozawa, Lei Wang, Matthias Troyer, and Yoshiro Takahashi (2016), “Topological Thouless pumping of ultracold fermions,” Nat. Phys. 12 (4), 296–300.
- Nandkishore and Hermele [2019] Nandkishore, Rahul M, and Michael Hermele (2019), “Fractons,” Annu. Rev. Condens. Matter Phys. 10 (1), 295–313.
- Nayak et al. [2008] Nayak, Chetan, Steven H. Simon, Ady Stern, Michael Freedman, and Sankar Das Sarma (2008), “Non-Abelian anyons and topological quantum computation,” Rev. Mod. Phys. 80, 1083–1159.
- Newkome et al. [2006] Newkome, George R, Pingshan Wang, Charles N. Moorefield, Tae Joon Cho, Prabhu P. Mohapatra, Sinan Li, Seok-Ho Hwang, Olena Lukoyanova, Luis Echegoyen, Judith A. Palagallo, Violeta Iancu, and Saw-Wai Hla (2006), “Nanoassembly of a Fractal Polymer: A Molecular ”Sierpinski Hexagonal Gasket”,” Science 312 (5781), 1782–1785.
- Brzezińska et al. [2018] Brzezińska, Marta, Ashley M. Cook, and Titus Neupert (2018), “Topology in the Sierpiński-Hofstadter problem,” Phys. Rev. B 98, 205116.
- Nguyen et al. [2023] Nguyen, D-H-Minh, Chiara Devescovi, Dung Xuan Nguyen, Hai Son Nguyen, and Dario Bercioux (2023), “Fermi arc reconstruction in synthetic photonic lattice,” Phys. Rev. Lett. 131, 053602.
- Nguyen et al. [2021] Nguyen, H C, Dung Xuan Nguyen, Thibaud Louvet, Xavier Letartre, Pierre Viktorovitch, and Hai Son Nguyen (2021), “Topological properties of photonic bands with synthetic momentum,” arXiv 10.48550/arxiv.2111.02843, 2111.02843 .
- Nguyen et al. [2015] Nguyen, H S, D. Gerace, I. Carusotto, D. Sanvitto, E. Galopin, A. Lemaître, I. Sagnes, J. Bloch, and A. Amo (2015), “Acoustic black hole in a stationary hydrodynamic flow of microcavity polaritons,” Phys. Rev. Lett. 114, 036402.
- Muñoz de Nova et al. [2019] Muñoz de Nova, Juan Ramón, Katrine Golubkov, Victor I. Kolobov, and Jeff Steinhauer (2019), “Observation of thermal Hawking radiation and its temperature in an analogue black hole,” Nature 569 (7758), 688–691.
- Ohtsubo et al. [2022] Ohtsubo, Yoshiyuki, Takuto Nakamura, and Shin-ichi Kimura (2022), “Growth of 2D topological material Bi on InSb(111)B with fractal surface structures,” arXiv 10.48550/arxiv.2212.14493, 2212.14493 .
- Oka and Kitamura [2019] Oka, Takashi, and Sota Kitamura (2019), “Floquet engineering of quantum materials,” Annu. Rev. Condens. Matter Phys. 10 (1), 387–408.
- Ostlund et al. [1983] Ostlund, Stellan, Rahul Pandit, David Rand, Hans Joachim Schellnhuber, and Eric D. Siggia (1983), “One-Dimensional Schrödinger Equation with an Almost Periodic Potential,” Phys. Rev. Lett. 50, 1873–1876.
- Ozawa et al. [2019] Ozawa, Tomoki, Hannah M. Price, Alberto Amo, Nathan Goldman, Mohammad Hafezi, Ling Lu, Mikael C. Rechtsman, David Schuster, Jonathan Simon, Oded Zilberberg, and Iacopo Carusotto (2019), “Topological photonics,” Rev. Mod. Phys. 91, 015006.
- Ozawa et al. [2016] Ozawa, Tomoki, Hannah M. Price, Nathan Goldman, Oded Zilberberg, and Iacopo Carusotto (2016), “Synthetic dimensions in integrated photonics: From optical isolation to four-dimensional quantum Hall physics,” Phys. Rev. A 93, 043827.
- Pai and Prem [2019] Pai, Shriya, and Abhinav Prem (2019), “Topological states on fractal lattices,” Phys. Rev. B 100 (15), 155135.
- Pal and Huse [2010] Pal, Arijeet, and David A. Huse (2010), “Many-body localization phase transition,” Phys. Rev. B 82, 174411.
- Pellerin et al. [2024] Pellerin, F, R. Houvenaghel, W. A. Coish, I. Carusotto, and P. St-Jean (2024), “Wave-function tomography of topological dimer chains with long-range couplings,” Phys. Rev. Lett. 132, 183802.
- Penrose [1974] Penrose, R (1974), “The role of aesthetics in pure and applied mathematical research,” Bull.Inst.Math.Appl. 10, 266–271.
- Peri et al. [2019] Peri, Valerio, Marc Serra-Garcia, Roni Ilan, and Sebastian D. Huber (2019), “Axial-field-induced chiral channels in an acoustic Weyl system,” Nat. Phys. 15 (4), 357.
- Pietracaprina et al. [2018] Pietracaprina, Francesca, Nicolas Macé, David J. Luitz, and Fabien Alet (2018), “Shift-invert diagonalization of large many-body localizing spin chains,” SciPost Phys. 5, 045.
- Piquero-Zulaica et al. [2022] Piquero-Zulaica, Ignacio, Jorge Lobo-Checa, Zakaria M. Abd El-Fattah, J. Enrique Ortega, Florian Klappenberger, Willi Auwärter, and Johannes V. Barth (2022), “Engineering quantum states and electronic landscapes through surface molecular nanoarchitectures,” Rev. Mod. Phys. 94 (4), 045008.
- Poddubny et al. [2008] Poddubny, A N, L. Pilozzi, M. M. Voronov, and E. L. Ivchenko (2008), “Resonant Fibonacci quantum well structures in one dimension,” Phys. Rev. B 77, 113306.
- Pretko et al. [2020] Pretko, Michael, Xie Chen, and Yizhi You (2020), “Fracton phases of matter,” Int. J. Mod. Phys. A 35 (06), 2030003.
- Price [2022] Price, Hannah (2022), “Simulating four-dimensional physics in the laboratory,” Phys. Today 75 (4), 38–44.
- Price et al. [2017] Price, Hannah M, Tomoki Ozawa, and Nathan Goldman (2017), “Synthetic dimensions for cold atoms from shaking a harmonic trap,” Phys. Rev. A 95, 023607.
- Qi et al. [2008] Qi, Xiao-Liang, Taylor L. Hughes, and Shou-Cheng Zhang (2008), “Topological field theory of time-reversal invariant insulators,” Phys. Rev. B 78, 195424.
- Qi and Zhang [2011] Qi, Xiao-Liang, and Shou-Cheng Zhang (2011), “Topological insulators and superconductors,” Rev. Mod. Phys. 83, 1057–1110.
- Rafi-Ul-Islam et al. [2020] Rafi-Ul-Islam, S M, Zhuo Bin Siu, and Mansoor B. A. Jalil (2020), “Topoelectrical circuit realization of a Weyl semimetal heterojunction,” Commun. Phys. 3 (1), 72.
- Rajagopal et al. [2019] Rajagopal, Shankari V, Toshihiko Shimasaki, Peter Dotti, Mantas Račiūnas, Ruwan Senaratne, Egidijus Anisimovas, André Eckardt, and David M. Weld (2019), “Phasonic spectroscopy of a quantum gas in a quasicrystalline lattice,” Phys. Rev. Lett. 123, 223201.
- Rechtsman et al. [2013] Rechtsman, Mikael C, Julia M Zeuner, Yonatan Plotnik, Yaakov Lumer, Daniel Podolsky, Felix Dreisow, Stefan Nolte, Mordechai Segev, and Alexander Szameit (2013), “Photonic Floquet topological insulators,” Nature 496 (7444), 196–200.
- Rispoli et al. [2019] Rispoli, Matthew, Alexander Lukin, Robert Schittko, Sooshin Kim, M. Eric Tai, Julian Léonard, and Markus Greiner (2019), “Quantum critical behaviour at the many-body localization transition,” Nature 573 (7774), 385–389.
- Roati et al. [2008] Roati, Giacomo, Chiara D’Errico, Leonardo Fallani, Marco Fattori, Chiara Fort, Matteo Zaccanti, Giovanni Modugno, Michele Modugno, and Massimo Inguscio (2008), “Anderson localization of a non-interacting Bose–Einstein condensate,” Nature 453 (7197), 895–898.
- Rodríguez-Laguna et al. [2017] Rodríguez-Laguna, Javier, Leticia Tarruell, Maciej Lewenstein, and Alessio Celi (2017), “Synthetic Unruh effect in cold atoms,” Phys. Rev. A 95, 013627.
- Rojo-Francàs et al. [2024] Rojo-Francàs, Abel, Priyanshu Pansari, Utso Bhattacharya, Bruno Juliá-Díaz, and Tobias Grass (2024), “Anomalous quantum transport in fractal lattices,” arXiv:2401.16077 .
- Roushan et al. [2017] Roushan, P, C. Neill, J. Tangpanitanon, V. M. Bastidas, A. Megrant, R. Barends, Y. Chen, Z. Chen, B. Chiaro, A. Dunsworth, A. Fowler, B. Foxen, M. Giustina, E. Jeffrey, J. Kelly, E. Lucero, J. Mutus, M. Neeley, C. Quintana, D. Sank, A. Vainsencher, J. Wenner, T. White, H. Neven, D. G. Angelakis, and J. Martinis (2017), “Spectroscopic signatures of localization with interacting photons in superconducting qubits,” Science 358 (6367), 1175–1179.
- Roy et al. [2021] Roy, Shilpi, Tapan Mishra, B. Tanatar, and Saurabh Basu (2021), “Reentrant localization transition in a quasiperiodic chain,” Phys. Rev. Lett. 126, 106803.
- Rudner and Lindner [2020] Rudner, Mark S, and Netanel H. Lindner (2020), “Band structure engineering and non-equilibrium dynamics in Floquet topological insulators,” Nat. Rev. Phys. 2 (5), 229–244.
- Rudner et al. [2013] Rudner, Mark S, Netanel H. Lindner, Erez Berg, and Michael Levin (2013), “Anomalous edge states and the bulk-edge correspondence for periodically driven two-dimensional systems,” Phys. Rev. X 3, 031005.
- Ryu et al. [2010] Ryu, Shinsei, Andreas P Schnyder, Akira Furusaki, and Andreas W W Ludwig (2010), “Topological insulators and superconductors: tenfold way and dimensional hierarchy,” New J. Phys. 12 (6), 065010.
- Sabsovich et al. [2022] Sabsovich, Daniel, Paul Wunderlich, Victor Fleurov, Dmitry I. Pikulin, Roni Ilan, and Tobias Meng (2022), “Hawking fragmentation and Hawking attenuation in Weyl semimetals,” Phys. Rev. Res. 4, 013055.
- Sala et al. [2015] Sala, V G, D. D. Solnyshkov, I. Carusotto, T. Jacqmin, A. Lemaître, H. Terças, A. Nalitov, M. Abbarchi, E. Galopin, I. Sagnes, J. Bloch, G. Malpuech, and A. Amo (2015), “Spin-orbit coupling for photons and polaritons in microstructures,” Phys. Rev. X 5, 011034.
- Salamon et al. [2020] Salamon, Tymoteusz, Alessio Celi, Ravindra W. Chhajlany, Irénée Frérot, Maciej Lewenstein, Leticia Tarruell, and Debraj Rakshit (2020), “Simulating twistronics without a twist,” Phys. Rev. Lett. 125, 030504.
- Sanchez-Palencia and Santos [2005] Sanchez-Palencia, L, and L. Santos (2005), “Bose-Einstein condensates in optical quasicrystal lattices,” Phys. Rev. A 72, 053607.
- Sbroscia et al. [2020] Sbroscia, Matteo, Konrad Viebahn, Edward Carter, Jr-Chiun Yu, Alexander Gaunt, and Ulrich Schneider (2020), “Observing Localization in a 2D Quasicrystalline Optical Lattice,” Phys. Rev. Lett. 125, 200604.
- Schlosser et al. [2001] Schlosser, Nicolas, Georges Reymond, Igor Protsenko, and Philippe Grangier (2001), “Sub-poissonian loading of single atoms in a microscopic dipole trap,” Nature 411 (6841), 1024–1027.
- Schmidt and Koch [2013] Schmidt, Sebastian, and Jens Koch (2013), “Circuit QED lattices: Towards quantum simulation with superconducting circuits,” Ann. Phys. (Berl.) 525 (6), 395–412.
- Schneider et al. [2022] Schneider, Lucas, Philip Beck, Jannis Neuhaus-Steinmetz, Levente Rózsa, Thore Posske, Jens Wiebe, and Roland Wiesendanger (2022), “Precursors of Majorana modes and their length-dependent energy oscillations probed at both ends of atomic Shiba chains,” Nat. Nanotechnol. 17 (4), 384–389.
- Schreiber et al. [2015] Schreiber, Michael, Sean S. Hodgman, Pranjal Bordia, Henrik P. Lüschen, Mark H. Fischer, Ronen Vosk, Ehud Altman, Ulrich Schneider, and Immanuel Bloch (2015), “Observation of many-body localization of interacting fermions in a quasirandom optical lattice,” Science 349 (6250), 842–845.
- Schwartz et al. [2007] Schwartz, Tal, Guy Bartal, Shmuel Fishman, and Mordechai Segev (2007), “Transport and Anderson localization in disordered two-dimensional photonic lattices,” Nature 446 (7131), 52–55.
- Semeghini et al. [2021] Semeghini, G, H. Levine, A. Keesling, S. Ebadi, T. T. Wang, D. Bluvstein, R. Verresen, H. Pichler, M. Kalinowski, R. Samajdar, A. Omran, S. Sachdev, A. Vishwanath, M. Greiner, V. Vuletic, and M. D. Lukin (2021), “Probing topological spin liquids on a programmable quantum simulator,” Science 374 (6572), 1242–1247.
- Setiawan et al. [2017] Setiawan, F, Dong-Ling Deng, and J. H. Pixley (2017), “Transport properties across the many-body localization transition in quasiperiodic and random systems,” Phys. Rev. B 96, 104205.
- Shang et al. [2015] Shang, Jian, Yongfeng Wang, Min Chen, Jingxin Dai, Xiong Zhou, Julian Kuttner, Gerhard Hilt, Xiang Shao, J. Michael Gottfried, and Kai Wu (2015), “Assembling molecular Sierpiński triangle fractals,” Nat. Chem. 7 (5), 389–393.
- Shechtman et al. [1984] Shechtman, D, I. Blech, D. Gratias, and J. W. Cahn (1984), “Metallic phase with long-range orientational order and no translational symmetry,” Phys. Rev. Lett. 53, 1951–1953.
- Sheng et al. [2022] Sheng, Chong, Yao Wang, Yijun Chang, Huiming Wang, Yongheng Lu, Yingyue Yang, Shining Zhu, Xianmin Jin, and Hui Liu (2022), “Bound vortex light in an emulated topological defect in photonic lattices,” Light Sci. Appl. 11 (1), 243.
- Shockley [1939] Shockley, William (1939), “On the surface states associated with a periodic potential,” Phys. Rev. 56, 317–323.
- Sierant et al. [2024] Sierant, Piotr, Maciej Lewenstein, Antonello Scardicchio, Lev Vidmar, and Jakub Zakrzewski (2024), “Many-body localization in the age of classical computing,” arXiv:2403.07111 [cond-mat.dis-nn] .
- Sierant et al. [2020] Sierant, Piotr, Maciej Lewenstein, and Jakub Zakrzewski (2020), “Polynomially filtered exact diagonalization approach to many-body localization,” Phys. Rev. Lett. 125, 156601.
- Slot et al. [2019] Slot, M R, S. N. Kempkes, E. J. Knol, W. M. J. van Weerdenburg, J. J. van den Broeke, D. Wegner, D. Vanmaekelbergh, A. A. Khajetoorians, C. Morais Smith, and I. Swart (2019), “-Band Engineering in Artificial Electronic Lattices,” Phys. Rev. X 9, 011009.
- Slot et al. [2017] Slot, Marlou R, Thomas S Gardenier, Peter H Jacobse, Guido C P van Miert, Sander N Kempkes, Stephan J M Zevenhuizen, Cristiane Morais Smith, Daniel Vanmaekelbergh, and Ingmar Swart (2017), “Experimental realization and characterization of an electronic Lieb lattice,” Nat. Phys. 13 (7), 672 – 676.
- Smerdon et al. [2014] Smerdon, J A, K. M. Young, M. Lowe, S. S. Hars, T. P. Yadav, D. Hesp, V. R. Dhanak, A. P. Tsai, H. R. Sharma, and R. McGrath (2014), “Templated Quasicrystalline Molecular Ordering,” Nano Lett. 14 (3), 1184–1189.
- Soldini et al. [2023] Soldini, Martina O, Felix Küster, Glenn Wagner, Souvik Das, Amal Aldarawsheh, Ronny Thomale, Samir Lounis, Stuart S. P. Parkin, Paolo Sessi, and Titus Neupert (2023), “Two-dimensional Shiba lattices as a possible platform for crystalline topological superconductivity,” Nat. Phys. 19 (12), 1848–1854.
- Soukoulis and Economou [1982] Soukoulis, C M, and E. N. Economou (1982), “Localization in one-dimensional lattices in the presence of incommensurate potentials,” Phys. Rev. Lett. 48, 1043–1046.
- Spar et al. [2022] Spar, Benjamin M, Elmer Guardado-Sanchez, Sungjae Chi, Zoe Z. Yan, and Waseem S. Bakr (2022), “Realization of a Fermi-Hubbard Optical Tweezer Array,” Phys. Rev. Lett. 128, 223202.
- Spurrier and Cooper [2020] Spurrier, Stephen, and Nigel R. Cooper (2020), “Kane-mele with a twist: Quasicrystalline higher-order topological insulators with fractional mass kinks,” Phys. Rev. Res. 2, 033071.
- Stampfli [1986] Stampfli, P (1986), “A dodecagonal quasiperiodic lattice in two dimensions.” Helv. Phys. Acta 59, 1260–1263.
- Steinhauer [2016] Steinhauer, Jeff (2016), “Observation of quantum Hawking radiation and its entanglement in an analogue black hole,” Nat. Phys. 12 (10), 959–965.
- Stuhl et al. [2015] Stuhl, B K, H.-I. Lu, L. M. Aycock, D. Genkina, and I. B. Spielman (2015), “Visualizing edge states with an atomic Bose gas in the quantum Hall regime,” Science 349 (6255), 1514–1518.
- Stålhammar et al. [2023] Stålhammar, Marcus, Jorge Larana-Aragon, Lukas Rødland, and Flore K Kunst (2023), “ symmetry-protected exceptional cones and analogue Hawking radiation,” New J. Phys. 25 (4), 043012.
- Su et al. [1979] Su, W P, J. R. Schrieffer, and A. J. Heeger (1979), “Solitons in polyacetylene,” Phys. Rev. Lett. 42, 1698–1701.
- Szabó and Schneider [2020] Szabó, Attila, and Ulrich Schneider (2020), “Mixed spectra and partially extended states in a two-dimensional quasiperiodic model,” Phys. Rev. B 101, 014205.
- Tanese et al. [2014] Tanese, D, E. Gurevich, F. Baboux, T. Jacqmin, A. Lemaître, E. Galopin, I. Sagnes, A. Amo, J. Bloch, and E. Akkermans (2014), “Fractal Energy Spectrum of a Polariton Gas in a Fibonacci Quasiperiodic Potential,” Phys. Rev. Lett. 112, 146404.
- Tang and Kohmoto [1986] Tang, Chao, and Mahito Kohmoto (1986), “Global scaling properties of the spectrum for a quasiperiodic Schrödinger equation,” Phys. Rev. B 34, 2041–2044.
- Tang et al. [2018] Tang, Hao, Xiao-Feng Lin, Zhen Feng, Jing-Yuan Chen, Jun Gao, Ke Sun, Chao-Yue Wang, Peng-Cheng Lai, Xiao-Yun Xu, Yao Wang, Lu-Feng Qiao, Ai-Lin Yang, and Xian-Min Jin (2018), “Experimental two-dimensional quantum walk on a photonic chip,” Sci. Adv. 4 (5), eaat3174.
- Tang et al. [2023a] Tang, Haoning, Beicheng Lou, Fan Du, Guangqi Gao, Mingjie Zhang, Xueqi Ni, Evelyn Hu, Amir Yacoby, Yuan Cao, Shanhui Fan, and Eric Mazur (2023a), “On-chip multidimensional dynamic control of twisted moiré photonic crystal for smart sensing and imaging,” arXiv:2312.09089 [physics.optics] .
- Tang et al. [2023b] Tang, Haoning, Beicheng Lou, Fan Du, Mingjie Zhang, Xueqi Ni, Weijie Xu, Rebekah Jin, Shanhui Fan, and Eric Mazur (2023b), “Experimental probe of twist angle–dependent band structure of on-chip optical bilayer photonic crystal,” Sci. Adv. 9 (28), eadh8498.
- Tao and Xu [2023] Tao, Yu-Liang, and Yong Xu (2023), “Higher-order topological hyperbolic lattices,” Phys. Rev. B 107, 184201.
- Tarquini et al. [2017] Tarquini, E, G. Biroli, and M. Tarzia (2017), “Critical properties of the Anderson localization transition and the high-dimensional limit,” Phys. Rev. B 95, 094204.
- Thouless [1983] Thouless, D J (1983), “Quantization of particle transport,” Phys. Rev. B 27, 6083–6087.
- Thouless et al. [1982] Thouless, D J, M. Kohmoto, M. P. Nightingale, and M. den Nijs (1982), “Quantized Hall Conductance in a Two-Dimensional Periodic Potential,” Phys. Rev. Lett. 49, 405–408.
- Tian et al. [2023] Tian, Weikun, Wen Jun Wee, An Qu, Billy Jun Ming Lim, Prithvi Raj Datla, Vanessa Pei Wen Koh, and Huanqian Loh (2023), “Parallel assembly of arbitrary defect-free atom arrays with a multitweezer algorithm,” Phys. Rev. Appl. 19, 034048.
- Todd et al. [1986] Todd, J, R. Merlin, Roy Clarke, K. M. Mohanty, and J. D. Axe (1986), “Synchrotron X-Ray Study of a Fibonacci Superlattice,” Phys. Rev. Lett. 57, 1157–1160.
- Tononi et al. [2024] Tononi, A, G. E. Astrakharchik, and D. S. Petrov (2024), “Gas-to-soliton transition of attractive bosons on a spherical surface,” AVS Quantum Sci. 6 (2), 023201.
- Tononi et al. [2022] Tononi, Andrea, Axel Pelster, and Luca Salasnich (2022), “Topological superfluid transition in bubble-trapped condensates,” Phys. Rev. Res. 4, 013122.
- Tononi and Salasnich [2024] Tononi, Andrea, and Luca Salasnich (2024), “Shell-shaped atomic gases,” Phys. Rep. 1072, 1–48.
- Unruh [1976] Unruh, W G (1976), “Notes on black-hole evaporation,” Phys. Rev. D 14, 870–892.
- Unruh [1981] Unruh, W G (1981), “Experimental black-hole evaporation?” Phys. Rev. Lett. 46, 1351–1353.
- Uri et al. [2023] Uri, Aviram, Sergio C. de la Barrera, Mallika T. Randeria, Daniel Rodan-Legrain, Trithep Devakul, Philip J. D. Crowley, Nisarga Paul, Kenji Watanabe, Takashi Taniguchi, Ron Lifshitz, Liang Fu, Raymond C. Ashoori, and Pablo Jarillo-Herrero (2023), “Superconductivity and strong interactions in a tunable moiré quasicrystal,” Nature 620 (7975), 762–767.
- Urwyler et al. [2022] Urwyler, David M, Patrick M. Lenggenhager, Igor Boettcher, Ronny Thomale, Titus Neupert, and Tomá š Bzdušek (2022), “Hyperbolic topological band insulators,” Phys. Rev. Lett. 129, 246402.
- Varjas et al. [2019] Varjas, Dániel, Alexander Lau, Kim Pöyhönen, Anton R. Akhmerov, Dmitry I. Pikulin, and Ion Cosma Fulga (2019), “Topological phases without crystalline counterparts,” Phys. Rev. Lett. 123, 196401.
- Varma and Žnidarič [2019] Varma, Vipin Kerala, and Marko Žnidarič (2019), “Diffusive transport in a quasiperiodic Fibonacci chain: Absence of many-body localization at weak interactions,” Phys. Rev. B 100, 085105.
- Verbin et al. [2013] Verbin, Mor, Oded Zilberberg, Yaacov E. Kraus, Yoav Lahini, and Yaron Silberberg (2013), “Observation of topological phase transitions in photonic quasicrystals,” Phys. Rev. Lett. 110, 076403.
- Viebahn et al. [2019] Viebahn, Konrad, Matteo Sbroscia, Edward Carter, Jr-Chiun Yu, and Ulrich Schneider (2019), “Matter-wave diffraction from a quasicrystalline optical lattice,” Phys. Rev. Lett. 122, 110404.
- Vieira [2005] Vieira, André P (2005), “Low-energy properties of aperiodic quantum spin chains,” Phys. Rev. Lett. 94, 077201.
- Viermann et al. [2022] Viermann, Celia, Marius Sparn, Nikolas Liebster, Maurus Hans, Elinor Kath, Álvaro Parra-López, Mireia Tolosa-Simeón, Natalia Sánchez-Kuntz, Tobias Haas, Helmut Strobel, Stefan Floerchinger, and Markus K. Oberthaler (2022), “Quantum field simulator for dynamics in curved spacetime,” Nature 611 (7935), 260–264.
- Vijay et al. [2016] Vijay, Sagar, Jeongwan Haah, and Liang Fu (2016), “Fracton topological order, generalized lattice gauge theory, and duality,” Phys. Rev. B 94, 235157.
- Volovik [2016] Volovik, G E (2016), “Black hole and Hawking radiation by type-II Weyl fermions,” JETP Lett. 104 (9), 645–648.
- Volovik and Zhang [2017] Volovik, G E, and K. Zhang (2017), “Lifshitz Transitions, Type-II Dirac and Weyl Fermions, Event Horizon and All That,” J. Low Temp. Phys. 189 (5-6), 276–299.
- Štrkalj et al. [2021] Štrkalj, Antonio, Elmer V. H. Doggen, Igor V. Gornyi, and Oded Zilberberg (2021), “Many-body localization in the interpolating Aubry-André-Fibonacci model,” Phys. Rev. Res. 3, 033257.
- Wang et al. [2018] Wang, Bo, Tian Chen, and Xiangdong Zhang (2018), “Experimental Observation of Topologically Protected Bound States with Vanishing Chern Numbers in a Two-Dimensional Quantum Walk,” Phys. Rev. Lett. 121, 100501.
- Wang et al. [2016] Wang, Jun, Xia-Ji Liu, Gao Xianlong, and Hui Hu (2016), “Phase diagram of a non-Abelian Aubry-André-Harper model with -wave superfluidity,” Phys. Rev. B 93, 104504.
- Wang et al. [2021] Wang, Kai, Avik Dutt, Charles C. Wojcik, and Shanhui Fan (2021), “Topological complex-energy braiding of non-Hermitian bands,” Nature 598 (7879), 59–64.
- Wang [1995] Wang, Xiang Rong (1995), “Localization in fractal spaces: Exact results on the Sierpinski gasket,” Phys. Rev. B 51, 9310–9313.
- Wang et al. [2020a] Wang, Yibo, Sayali Shevate, Tobias Martin Wintermantel, Manuel Morgado, Graham Lochead, and Shannon Whitlock (2020a), “Preparation of hundreds of microscopic atomic ensembles in optical tweezer arrays,” npj Quantum Inf. 6 (1), 1–5.
- Wang et al. [2020b] Wang, You, Hannah M. Price, Baile Zhang, and Y. D. Chong (2020b), “Circuit implementation of a four-dimensional topological insulator,” Nat. Commun. 11 (1), 2356.
- Wang et al. [2022] Wang, Yunfei, Jia-Hui Zhang, Yuqing Li, Jizhou Wu, Wenliang Liu, Feng Mei, Ying Hu, Liantuan Xiao, Jie Ma, Cheng Chin, and Suotang Jia (2022), “Observation of Interaction-Induced Mobility Edge in an Atomic Aubry-André Wire,” Phys. Rev. Lett. 129, 103401.
- Watanabe et al. [2018] Watanabe, Haruki, Hoi Chun Po, and Ashvin Vishwanath (2018), “Structure and topology of band structures in the 1651 magnetic space groups,” Sci. Adv. 4 (8), eaat8685.
- Weidemann et al. [2020] Weidemann, Sebastian, Mark Kremer, Tobias Helbig, Tobias Hofmann, Alexander Stegmaier, Martin Greiter, Ronny Thomale, and Alexander Szameit (2020), “Topological funneling of light,” Science 368 (6488), 311–314.
- Weidinger et al. [2018] Weidinger, Simon A, Sarang Gopalakrishnan, and Michael Knap (2018), “Self-consistent Hartree-Fock approach to many-body localization,” Phys. Rev. B 98, 224205.
- Weitenberg and Simonet [2021] Weitenberg, Christof, and Juliette Simonet (2021), “Tailoring quantum gases by Floquet engineering,” Nat. Phys. 17 (12), 1342–1348, 2102.07009 .
- Werchner et al. [2009] Werchner, M, M. Schafer, M. Kira, S. W. Koch, J. Sweet, J. D. Olitzky, J. Hendrickson, B. C. Richards, G. Khitrova, H. M. Gibbs, A. N. Poddubny, E. L. Ivchenko, M. Voronov, and M. Wegener (2009), “One dimensional resonant Fibonacci quasicrystals: noncanonical linear and canonical nonlinear effects,” Opt. Express 17 (8), 6813–6828.
- Wilk et al. [2010] Wilk, T, A. Gaëtan, C. Evellin, J. Wolters, Y. Miroshnychenko, P. Grangier, and A. Browaeys (2010), “Entanglement of Two Individual Neutral Atoms Using Rydberg Blockade,” Phys. Rev. Lett. 104, 010502.
- Wimmer et al. [2017] Wimmer, Martin, Hannah M. Price, Iacopo Carusotto, and Ulf Peschel (2017), “Experimental measurement of the Berry curvature from anomalous transport,” Nat. Phys. 13 (6), 545–550.
- Wimmer et al. [2015] Wimmer, Martin, Alois Regensburger, Mohammad-Ali Miri, Christoph Bersch, Demetrios N. Christodoulides, and Ulf Peschel (2015), “Observation of optical solitons in -symmetric lattices,” Nat. Commun. 6 (1), 7782.
- Windpassinger and Sengstock [2013] Windpassinger, Patrick, and Klaus Sengstock (2013), “Engineering novel optical lattices,” Rep. Prog. Phys. 76 (8), 086401.
- Witten [1998] Witten, Edward (1998), “Anti de Sitter space and holography,” Adv. Theor. Math. Phys. 2 (2), 253–291.
- Xu et al. [2019] Xu, Shenglong, Xiao Li, Yi-Ting Hsu, Brian Swingle, and S. Das Sarma (2019), “Butterfly effect in interacting Aubry-Andre model: Thermalization, slow scrambling, and many-body localization,” Phys. Rev. Res. 1, 032039.
- Xu et al. [2021] Xu, Xiao-Yun, Xiao-Wei Wang, Dan-Yang Chen, C. Morais Smith, and Xian-Min Jin (2021), “Quantum transport in fractal networks,” Nat. Photonics 15 (9), 703–710.
- Xu and Duan [2016] Xu, Yong, and L.-M. Duan (2016), “Type-II Weyl points in three-dimensional cold-atom optical lattices,” Phys. Rev. A 94, 053619.
- Yahyavi et al. [2019] Yahyavi, M, B. Hetényi, and B. Tanatar (2019), “Generalized Aubry-André-Harper model with modulated hopping and -wave pairing,” Phys. Rev. B 100, 064202.
- Yang and Zhang [2016] Yang, Zhaoju, and Baile Zhang (2016), “Acoustic Type-II Weyl Nodes from Stacking Dimerized Chains,” Phys. Rev. Lett. 117, 224301.
- Yao and Zakrzewski [2020] Yao, Ruixiao, and Jakub Zakrzewski (2020), “Many-body localization in the Bose-Hubbard model: Evidence for mobility edge,” Phys. Rev. B 102, 014310.
- Yao et al. [2018] Yao, Wei, Eryin Wang, Changhua Bao, Yiou Zhang, Kenan Zhang, Kejie Bao, Chun Kai Chan, Chaoyu Chen, Jose Avila, Maria C. Asensio, Junyi Zhu, and Shuyun Zhou (2018), “Quasicrystalline 30° twisted bilayer graphene as an incommensurate superlattice with strong interlayer coupling,” Proc. Natl. Acad. Sci. U.S.A. 115 (27), 6928–6933.
- Yazdani et al. [2023] Yazdani, Ali, Felix von Oppen, Bertrand I. Halperin, and Amir Yacoby (2023), “Hunting for Majoranas,” Science 380 (6651), eade0850.
- Yu et al. [2023] Yu, Jr-Chiun, Shaurya Bhave, Lee Reeve, Bo Song, and Ulrich Schneider (2023), “Observing the two-dimensional Bose glass in an optical quasicrystal,” arXiv:2303.00737 [cond-mat.quant-gas] .
- Yu et al. [2020] Yu, Rui, Y X Zhao, and Andreas P Schnyder (2020), “4D spinless topological insulator in a periodic electric circuit,” Natl. Sci. Rev. 7 (8), 1288–1295.
- Yuan et al. [2016] Yuan, Luqi, Yu Shi, and Shanhui Fan (2016), “Photonic gauge potential in a system with a synthetic frequency dimension,” Opt. Lett. 41 (4), 741.
- Zeng et al. [2016] Zeng, Qi-Bo, Shu Chen, and Rong Lü (2016), “Generalized Aubry-André-Harper model with -wave superconducting pairing,” Phys. Rev. B 94, 125408.
- Zhang et al. [2021] Zhang, Ren, Chenwei Lv, Yangqian Yan, and Qi Zhou (2021), “Efimov-like states and quantum funneling effects on synthetic hyperbolic surfaces,” Science Bulletin 66 (19), 1967–1972.
- Zhang et al. [2023] Zhang, Weixuan, Fengxiao Di, Xingen Zheng, Houjun Sun, and Xiangdong Zhang (2023), “Hyperbolic band topology with non-trivial second Chern numbers,” Nat. Commun. 14 (1), 1–9.
- Zhang et al. [2022] Zhang, Weixuan, Hao Yuan, Na Sun, Houjun Sun, and Xiangdong Zhang (2022), “Observation of novel topological states in hyperbolic lattices,” Nat. Commun. 13 (1), 2937, 2203.03214 .
- Zhao [2018] Zhao, Erhai (2018), “Topological circuits of inductors and capacitors,” Ann. Phys. (N. Y.) 399, 289–313.
- Zhong et al. [2020] Zhong, Han-Sen, Hui Wang, Yu-Hao Deng, Ming-Cheng Chen, Li-Chao Peng, Yi-Han Luo, Jian Qin, Dian Wu, Xing Ding, Yi Hu, Peng Hu, Xiao-Yan Yang, Wei-Jun Zhang, Hao Li, Yuxuan Li, Xiao Jiang, Lin Gan, Guangwen Yang, Lixing You, Zhen Wang, Li Li, Nai-Le Liu, Chao-Yang Lu, and Jian-Wei Pan (2020), “Quantum computational advantage using photons,” Science 370 (6523), 1460–1463.
- Zhu et al. [2023] Zhu, Zhaoxuan, Hepeng Yao, and Laurent Sanchez-Palencia (2023), “Thermodynamic Phase Diagram of Two-Dimensional Bosons in a Quasicrystal Potential,” Phys. Rev. Lett. 130, 220402.
- Zhu et al. [2024] Zhu, Zhaoxuan, Shengjie Yu, Dean Johnstone, and Laurent Sanchez-Palencia (2024), “Localization and spectral structure in two-dimensional quasicrystal potentials,” Phys. Rev. A 109, 013314.
- Zilberberg et al. [2018] Zilberberg, Oded, Sheng Huang, Jonathan Guglielmon, Mohan Wang, Kevin P. Chen, Yaacov E. Kraus, and Mikael C. Rechtsman (2018), “Photonic topological boundary pumping as a probe of 4D quantum Hall physics,” Nature 553 (7686), 59–62.